### v2

Abstract
We numerically examine dynamo action generated by a flow of an electrically conducting fluid in a precessing cylindrical cavity. We compare a simplified kinematic approach based on the solution of the magnetic induction equation with a prescribed velocity field with the results from a self-consistent three-dimensional simulation of the complete set of magnetohydrodynamic equations. The different kinematic models show a strong influence of the electromagnetic properties of outer layers, which manifests itself in the existence of two different branches with dynamo action. In all cases, we observe a minimum for the onset of dynamo action in a transitional regime, within which the hydrodynamic flow undergoes a change from a large-scale to a more small-scale, turbulent behaviour. However, significant differences in the absolute values for the critical magnetic Reynolds number occur, and they can be explained by the low rotation rate in the MHD simulations achievable for technical reasons. In these models the character of the dynamo solution is small-scale and the magnetic energy remains significantly smaller than the kinetic energy of the flow. In irregular intervals, we observe dynamo bursts with a local concentration of the magnetic field, resulting in a global increase of the magnetic energy by a factor of 3 to 5. However, diffusion of the local patches caused by strong local shear is too rapid, causing these features to exist for only a short period (i.e., shorter than a rotation period) so that their dynamical impact on the dynamo remains small. As the magnetic field is small-scale and weak, the nonlinear feedback on the flow through the Lorentz force remains small and arises essentially in terms of a slight damping of the fast timescales, whereas there is no noticeable change in flow amplitude compared to the hydrodynamic case.
I Introduction
The experimental investigation of magnetic field generation through a flow of electrically conductive fluid is of great interest, as such a dynamo process is crucial for many astrophysical bodies and allows conclusions to be drawn about their evolution and internal structure. Following the successful first-generation dynamo experiments in Karlsruhe [1] and Riga [2], which confirmed the fundamental principle, and subsequent dynamo experiments in Cadarache [3], characterized by a wide variety of different dynamic behaviours, a new experiment is currently being set up at HZDR [4, 5]. In this experiment, a flow of liquid sodium in a precessing cylinder is intended to generate a magnetic field. A mechanical drive such as precession is indeed capable of generating strong flows, allowing for the large flow amplitudes required for the occurrence of the dynamo effect. This contrasts with potentially more straightforward mechanisms, such as convection, which at typical laboratory scales may not achieve sufficiently vigorous flows due to lower driving efficiency. In addition, precession as a flow drive makes perfect sense in a dynamo experiment because driving on the largest possible length scale – the size of the container – opens up significantly more degrees of freedom for establishing a (turbulent) flow through cascade processes than in the case of the previously highly optimized experiments[6].
Precession is a repeatedly proposed driving mechanism for natural dynamos, whether for Earth [7] or the Moon [8]. Paleomagnetic records suggest that the geodynamo was active even before the formation of the Earth’s solid inner core, which implies that thermochemical convection, driven by the cooling and crystallization of the core, was not always the primary mechanism. Before the onset of thermochemical convection, alternative mechanisms must have driven the early geodynamo. Among the most plausible candidates are precession-driven and tide-driven flows, which could have provided the necessary energy for maintaining the magnetic field in the Earth’s early history. This follows from recent experiments and numerical computation yield rather large values for the heat conductivity under conditions typical for the Earth’s liquid core, which imposes strict energetic constraints for the convective state, particularly for the early geodynamo (before the formation of the solid inner Earth’s core, see e.g. Olson [9], Landeau et al. [10]). While it is relatively clear that the laminar flow directly driven by precession is energetically insignificant [11], the original idea of Malkus was based on a turbulent base state [7], which was underpinned by estimates of an upper limit to (viscous) dissipation by Kerswell [12]. Indeed, precession-driven fluid flows exhibit several instabilities, allowing both forward and backward cascades of energy transfer [13] that end up in a flow state with small-scale turbulence superimposed by large-scale flow contributions essentially in terms of a circulation opposite to the original rotation of the container [14, 15].
Regarding the Lunar dynamo, which around 4 billion years ago generated a magnetic field of comparable strength to that of today’s Earth, precession is indeed the most promising candidate to explain the observations including the disappearance of the dynamo about 3 billion years ago, see e.g. Stys and Dumberry [16].
Precession-driven dynamos have been the subject of numerous numerical studies, and there are now successful dynamo simulations in all conceivable geometries, including spheres [17], spheroids [18], cylinders [19], or cubes [20]. A preliminary experiment with sodium in a precessing cylinder was conducted in the early 1970s [21]. This experiment demonstrated the amplification of an externally applied magnetic field by a factor of three, however, the setup at that time was not large enough to achieve the magnetic Reynolds numbers required for a dynamo.
Preliminary investigations using a kinematic model based on the geometry of the planned experiment and the large-scale flow from hydrodynamic simulations showed that the dynamo effect works best in the narrow transition region that separates the subcritical from the supercritical state [4, 22]. In the present study, we deepen the investigation of the dynamo effect in the transitional regime and furthermore, we investigate the influence of the magnetic boundary conditions and turn to full self-consistent models to study the influence of the nonlinear terms and the Lorentz force feedback on the flow.
The present study is organized as follows. In section II.1, we introduce the equations that describe an electrically conducting fluid flow in a precessing cylinder and the related magnetic induction. We further summarize the essential response of the fluid to the forcing imposed by the precession of the container and present the evolution when increasing this forcing. The subsequent section II.3 is dedicated to the results obtained from kinematic dynamo models, where only the induction equation is solved numerically assuming a prescribed, time-independent flow. In this section, we detail the impact of the electrical conductivity of the container. Section III focuses on the corresponding results of the full set of magnetohydrodynamic (MHD) equations and reveals differences and similarities with the kinematic dynamo. We end our study with the conclusions in section IV, where we try to adopt our results for the forthcoming dynamo experiment at HZDR.
II Kinematic approach
II.1 Setup & Equations
Initially, we focus on the kinematic problem, where only the magnetic induction equation is solved and a time-independent velocity field is prescribed, which is supposed to represent the mean (time-averaged) velocity in the experiment. The main finding from this model is a value for the critical magnetic Reynolds number, which is related to the minimum amplitude of the velocity field required for a dynamo to occur. The kinematic approach does neither consider the back-reaction of the magnetic field on the fluid flow via the Lorentz force so that we have no possibility for any saturation, nor the impact of small-scale fluctuations that could hamper (turbulent diffusion) or support (-effect) the dynamo process.
We calculate the temporal evolution of the magnetic flux density (in the following called ’magnetic field’) by numerically solving the magnetic induction equation, which reads
In Eq. (1) represents the velocity field and is the magnetic Reynolds number defined with the angular velocity of the cylinder wit radius and the magnetic diffusivity , which is related to the electrical conductivity by (see also Figure 1a).
In all cases, the inner region of the simulation domain consists of a cylindrical container with radius and height (in dimensionless unities) which is filled with an electrically conducting fluid with a magnetic diffusivity . The fluid flow is prescribed by the data taken from hydrodynamic simulations presented in Pizzi, Giesecke, and Stefani [23], Pizzi et al. [13], Giesecke et al. [15]. In these studies, it is found that the flow in a cylinder emerging at a large nutation angle can be represented by the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow. A more detailed prescription of the structure of the flow field is given in the following subsection.
The flow fields used in this study are derived from hydrodynamic simulations with the spectral element code SEMTEX[24, 25]. A detailed description of the flow field as a function of the precession ratio defined by can be found in [23, 13, 14, 15]. In the following, we briefly summarize the main properties that are relevant for our present study. The flow in a cylinder precessing at a large nutation angle can be characterized by three regimes. In all cases, it is possible to represent the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow. The structure of the time-averaged flow fields as applied in the kinematic model is shown in Figure 2 for three precession ratios that are characteristic for the three flow regimes.
In the subcritical regime, the flow is essentially laminar, and the geometry can be described by the superposition of the rotational motion due to the rotating cylinder and the directly driven flow (essentially an mode proportional to , see left column in Figure 2). Overlaying this is the zonal flow maintained by the cylinder’s rotation (geostrophic and axisymmetric). Without precession, this would be a pure solid-body rotation with with but due to the influence of precession the radial profile of the zonal flow undergoes a significant modification which leads to a significant deceleration of the fluid rotation in the bulk (see bottom row in Figure 2). For larger forcing the flow transients into a supercritical region, in which the large-scale flow in the inner region has almost completely stopped so that the fluid essentially performs small-scale turbulent fluctuations and large-scale components are then only found near the walls of the container (see right column in Figure 2). These two regimes are linked by a transition zone, which is characterized by an intermittent behaviour of the large-scale flow with elements from both regimes. In this transitional region, there exists an additional large-scale contribution in the form of a double roll pattern as shown in the central column of Figure 2 (third row).
II.2 Code
II.2.1 Algorithm
Assuming a prescribed time-independent flow field, Eq. (1) specifies a linear problem and the solution follows with the complex eigenvalue consisting of the growth rate (real part ) and the frequency (imaginary part ). In order to solve Eq. (1) numerically, we use a finite volume approach in cylindrical coordinates. The scheme applies the constraint transport method to guarantee divergence-free solutions and is described in more detail in Giesecke, Stefani, and Gerbeth [26]. Extensive tests of the algorithm, in particular the impact of locally varying material properties, such as electrical conductivity and comparison with an alternative approach based on Spectral/Finite Elements (the SFEMaNS code) can be found in Giesecke et al. [27].
II.2.2 Boundary conditions
We apply pseudo-vacuum conditions for the magnetic field, which read (, where is the unit normal vector at the boundary so that the tangential components of the magnetic field an the outer boundary vanish. It is well known that pseudo-vacuum conditions in kinematic models usually underestimate the critical magnetic Reynolds number that must be exceeded for the dynamo to start. The reduction depends on the geometry, i.e. essentially on the aspect ratio, and amounted, for example, to up to 30% in the kinematic models of the VKS dynamo compared to insulator boundary conditions [27]. Nevertheless, we must restrict ourselves to the application of pseudo-vacuum conditions at the outer boundary of our system, because the strong shear near the outer boundaries enforces a high spatial resolution, which rules out the application of the Boundary-Element-Method (BEM, see e.g. Giesecke, Stefani, and Gerbeth [26]) for the consideration of insulating boundary conditions, as this method requires a large fully populated matrix to be kept in the memory. The size of this matrix scales with , where represents the number of grid points on the cylinder surface so that even current computers with a memory capacity of more than 1 TB are not sufficient to calculate the corresponding models. Instead, we estimate the impact of the magnetic boundary conditions employing various setups with one or two outer layers with thickness and as sketched in Figure 1. The underlying idea is that a sufficiently thick boundary layer with poor electrical conductivity (large magnetic diffusivity) provides a good shielding of the inner volume filled with the electrically conductive fluid so that the actual nature of the electrical boundary conditions sufficiently far away from the induction region is no longer important.
We examine three different geometric configurations. Initially we revise the results for a setup that considers only the fluid interior (i.e. ). The second case examines the impact of a container wall with different values of the magnetic diffusivity (i.e. ) and two different values for the thickness of the wall layer. In the thin layer model we use with a resolution of grid cells and in the thick layer model we use with a resolution of grid cells, whereby the magnetic diffusivity in the outer layer remains fixed at the maximum value of . Finally, in the fourth model we again apply a thin cylinder wall with and add a second outer (non-rotating) layer with and , which is supposed to resemble a non-conducting exteriour.
II.3 Results of the kinematic models
II.3.1 General overview
Previous kinematic models demonstrated that dynamo action is most effective when utilizing flow from the transitional region between the subcritical and supercritical states. A detailed investigation of which components of this realization are of decisive importance for the dynamo effect was carried out in Giesecke et al. [22] and showed that shear adjacent to the sidewall plays an important role as well as the presence of the double roll mode (DRM) which seems to be responsible for the considerable drop in the critical magnetic Reynolds number.
II.3.2 No outer layers
The typical behavior of the temporal evolution of the magnetic energy is shown in Figure 3a. Here we focus on two paradigmatic cases with the flow obtained at scaled to a magnetic Reynolds number of (red curve) and (blue curve), which is slightly below and above the dynamo threshold. A striking feature is the oscillating behavior of the energy, which is superimposed on the exponentially growing part. The oscillating character of the solution can be clearly seen in the animation, which can be found in the supplementary material to this study [28].
A systematic study of the impact of the velocity fields obtained at different the critical magnetic Reynolds number required for the onset of dynamo action which is shown in Figure 3b. We do not find any dynamos (at least up to ) for the flow fields in the laminar regime (i.e. below ). The lowest possible magnetic Reynolds number, is , which is obtained for the flow field from hydrodynamic simulations at . It is remarkable, that we find a significant smaller value for in the range than above, which coincides exactly with the transition regime of the velocity field from the subcritical to the supercritical state. We therefore assume that the axially symmetric double roll occurring exclusively in this parameter range plays a significant role for the onset of dynamo action at comparatively small magnetic Reynolds numbers.
II.3.3 Thin outer layer
In the following we focus on the velocity field that provides the minimum critical magnetic Reynolds number in our first models so that all results discussed for the kinematic approach in the reminder of this study are calculated from the velocity field obtained at (central column in Figure 2). We continue with the setup that considers an additional outer layer with thickness with increased magnetic diffusivity that emulates the container wall.
If we look at the temporal development of the magnetic energy and its spatial distribution, we recognize two different types of dynamo action. If the wall diffusivity is close to that of the fluid (i.e. ), a dynamo only occurs at very large (). This dynamo is time-independent (no oscillatory components, see orange and green curve in Figure 4a) and the associated field energy is essentially concentrated in a narrow region near the end caps (see iso-surfaces in Figure 4c). If the outer layer has an larger than (see blue curve in Figure 4b), we obtain an oscillating solution that already emerges at much smaller , which is exemplified by the blue curve for in Figure 4a. In that case the magnetic field energy is widely distributed throughout the entire volume (however still with a characteristic pattern, see Figure 4d.) The impact of the outer layer is illustrated by the growth rates shown in Figure 4b. When the magnetic diffusivity of the outer layer is increased to emulate the effect of a stainless steel container, an eigenmode is excited, the growth rates of which are represented by the inverted parabola in the individual curves in Figure 4b. This special shape results in positive growth rates for and for further increasing diffusivities the behaviour increasingly approaches that of pseudo-vaccum boundary conditions without outer layer (see black curve). It is therefore the eigenmode that belongs to the inverse parabolic growth rate, which corresponds to a dynamo that occurs at relatively low achievable in the experiment, and causes a magnetic field that permeates the entire volume. Interestingly, this eigenmode belongs to an oscillating dynamo while the much less efficient dynamo is stationary, although the opposite behavior is generally assumed [30].
II.3.4 Thick outer layer
The beneficial effect of the outer layer with sufficiently large diffusivity vanishes, when its thickness is increased, as can be recognized in Figure 5a which shows the evolution of the magnetic energy for while the magnetic diffusivity remains at . Here we end up with a critical magnetic Reynolds number slightly above and the solution is again characterized by a non-oscillatory behaviour with the magnetic energy constraint to two small section close to the endcaps, which clearly indicates that we are in the low efficiency branch (see Figure 5c). Obviously, the dominant solution in this case is again given by the less efficient branch.
Nevertheless, the highly efficient branch can be achieved again by assuming a more realistic configuration, where the container wall is emulated by a thin co-rotating layer with , while the laboratory is modeled by a thick outer layer with large diffusivity (actually, this should be an insulator in order to emulate air). In this case, we obtain an oscillating solution that occupies the entire cylinder volume and has a similar structure as in the case of the thin layer or pseudo-vacuum boundary conditions without an outer layer (see Figures 5b and 5d). However, the critical magnetic Reynolds number of is then about two and a half times as large as in these cases.
Thus we see from the kinematic linear models that there are two different branches of dynamo solutions, the realization of which depends on the configuration of the outer boundary regions, i.e. essentially on the boundary conditions. We have not investigated here at which exact parameters the transition from one branch to the other takes place, since the experimental realization must in any case start from a basically infinitely large outer volume of an insulator, namely the air in the laboratory. Interestingly, the resulting critical magnetic Reynolds number of is very close to the value derived by Goepfert and Tilgner [20] for the precession dynamo with the help of simulations in a cube.
In the following, we now want to solve the induction equation coupled to the Navier-Stokes equation with a comparable setup in a self-consistent MHD model, so that the time evolution of magnetic field and velocity field is solved simultaneously, including the feedback via the Lorentz force.
III Full set of magnetohydrodynamic equations
III.1 Code scheme
An extension of the kinematic problem arises when we additionally consider the temporal evolution of the flow field. In this case, the Navier-Stokes equation needs to be solved as well, which in the case of precession-driven flow takes the form
In addition to the terms commonly encountered in regular flow problems describing nonlinear interactions and dissipation, we have three contributions: the Coriolis force due to rotation, the Lorentz force due to the interaction with the magnetic field, and the Poincaré force accounting for acceleration due to perpetual change of the orientation of the cylinder’s axis of rotation. A detailed derivation of the Navier-Stokes equation for a precessing system is given e.g. in Tilgner [17].
We conduct direct numerical simulations (DNS) of the coupled system of equations (1) and (2) using the code SpecDyn[14, 31], which is based on a pseudo-spectral Fourier approach. The drawback of this approach is the requirement of periodic boundary conditions in all three directions. Since this rules out an application of a cylindrical geometry, the scheme is complemented by a direct forcing immersed boundary method (IBM) which allows the consideration of nearly arbitrary shaped boundaries that prescribe the confinement for the fluid flow when the boundary conditions on these boundaries are of no-slip type. The code is parallelized using a slab decomposition and a standard MPI library. Since the most important part of the calculations results from the computation of the Fast Fourier Transformation (FFT), special efforts are undertaken to apply an extremely efficient scheme The fulfilment of the divergence-free condition for the velocity field is ensured using a standard Pressure-Corrector method where the pressure serves as a Lagrangian multiplier so that is ensured (i.e. first, solve Equation (2) without pressure, then project the intermediate solution on the divergence-free space by calculating the necessary pressure field). The same scheme is also applied to the magnetic field to ensure that the solution of the induction equation is also divergence-free as required by Maxwell’s equations. The code has been tested with various configurations that proved accurate solutions with excellent convergence properties and nearly perfect strong scaling properties. A detailed description is given in Wilbert, Giesecke, and Grauer [14] and a comparable approach was used for numerical models of the VKS dynamo[32, 33].
III.2 Model set-up
We consider a set-up given by a cylindrical fluid region with radius and height embedded in a Cartesian domain with a edge length so that the size of the box provides sufficient space to let the magnetic field decline outside the cylindrical vessel. The cylindrical vessel is implemented by imposing internal conditions on the fluid that emulate the (virtual) container walls by enforcing no-slip conditions at the surface of the container applying the direct forcing immersed boundary method and we do not model any special region representing the wall of the vessel because of such a layer would not be properly resolved. Instead we use a magnetic diffusivity for the fluid region, whereas the ’non-liquid’ region outside the vessel has a magnetic diffusivity , which is supposed to emulate an electrically isolating exteriour. The simulations are carried out with a Reynolds number of which is the maximum possible at the resolution of that still sufficiently resolves the boundary layers at the fluid-wall interface of the vessel [31].
III.3 Results
Similar to the procedure in the kinematic case, the Poincaré number is varied, whereby different magnetic Reynolds numbers are considered, beginning with high values. If a positive mean growth rate is found for a particular set of parameters, further runs are executed consecutively lowering the value of until no more dynamo action can be observed. Rather similar to the kinematic models we find a limited range for the preferred occurrence of a dynamo in the interval . However, the critical magnetic Reynolds number is significantly higher with a minimum value of for . The slightly larger value for the optimum compared to the kinematic models can be explained by the (slightly) smaller (see also [15]). In the following we focus on the particular case and and compare the properties of this DNS solution with the kinematic approximation obtained at and . This paradigmatic examples well characterize the solution and allow a faster growth with shorter time period until a statistically steady state, while showing similar qualitative behavior as the dynamos found very close to the threshold. Here we focus on this particular and compare the properties of this DNS solution with and in the kinematic approximation.
III.3.1 Temporal evolution of the kinetic energy
The temporal evolutions of kinetic and magnetic energy are displayed in Fig. 6, where the kinetic and the magnetic energy are defined as
The simulations start with the pure hydrodynamic case and the effects by magnetic induction are switched on at . After a short initial phase the magnetic energy begins to grow for about cylinder rotations and finally goes into saturation. The induced magnetic energy is much smaller than and both energies are dominated by the contributions from the components. We find that, even in the saturated state, the presence of the magnetic field does not substantially alter the structure and the amplitude of the fluid flow via the Lorentz force. A striking property is that unlike the kinetic energy, which remains relatively steady, the magnetic energy is characterized by quasi-periodic peaks where its value suddenly rises to times above the average.
Although there are no substantial changes in the flow field due to the back-reaction of the magnetic field, the bursts can be clearly assigned to corresponding peaks in the interaction parameter , which compares the amplitude of the Lorentz force to the inertial term in the Navier-Stokes equation:
which in case of a dynamo follows Miralles, Plihon, and Pinton [34].
III.3.2 Dynamo bursts and field structure
The peaks in the magneitc energy are always associated with a very strong concentration of magnetic energy in the form of local patches, as can be seen in the three-dimensional representation (see Fig 8). These patches are short-lived despite the strong concentration of magnetic energy, and even if they appear regularly, no corresponding period can be assigned. The field structure obtained from the simulations is visualized in Figure 8 which shows nested isosurfaces of the magnetic energy density for a couple of typical snapshots. We find a small scale field in terms of seemingly randomly distributed in space and time. The magnetic field patches are slightly elongated in the azimuthal direction (which goes along with the decrease of the scale in the perpendicular direction) and in general we find a spatial scale of the order of the energy carrying turbulence. The magnetic field patches exhibit an intermittent character and the thickness of the structures at least in one dimension is determined by the resistive scale. Due to the special structure of the Poincare force and the associated mean flow in the form of a standing forced inertial wave, we are not dealing with an isotropic, homogeneous turbulence, so that Kazantsev’s theory (see e.g. [brandenburgXY]) is not directly applicable.
Let us label the second (major) peak as case A, the third peak as case B, and the non-peak solution as case C (see Fig. 6). The magnetic field is small-scale in all cases, whereas the velocity field is large-scale. Although our self-exciting dynamo is small-scale, ongoing hydromagnetic processes tend to excite larger scales (i.e., a large-scale magnetic field), as seen in Figure 8a. and 8b. However, these large-scale magnetic field structures are quite isolated; they do not extend throughout the entire volume and quickly dissipate due to strong magnetic diffusion, typically within 1-2 rotation periods.
In contrast, the velocity field is large-scale, characterized by two vortex structures near the cylinder wall, and is concentrated mainly near the cylinder boundary, while it is almost zero in the central volume due to the dominance of the component. The vorticity, , naturally follows the velocity characteristics, being concentrated at the cylinder’s border and negligible toward the central volume, as shown in Fig. LABEL:B-U-O. The velocity field remains largely unchanged in cases A and B, retaining characteristics similar to those in case C. The vorticity, however, penetrates slightly deeper into the central volume compared to case C (see Fig. LABEL:B-U-O), which represents the standard off-peak solution, where the magnetic field is weak and small-scale without any significant large-scale structures.
While the distribution of the magnetic energy in individual snapshots appears to be largely chaotic (with the exception of the bursts), this is no longer the case when we look at the time average in the turntable system. With sufficiently long averaging of more than 200 rotation periods, we get the distribution shown in Figure 9.
III.3.3 Backreaction and kinetic power spectrum
A reduction in the speed amplitude is also not absolutely necessary, because the dynamo that sets in is a small-scale dynamo. For this case, theoretical considerations have shown that (in a mean-field model) the feedback effect via the Lorentz force corresponds to an increased diffusivity or hyperdiffusivity [35]. However, an reaction of the fluid flow can be seen in a subtle way when looking at the temporal power-spectra of kinetic energy. Figure 10 compares the temporal Power-spectrum of the kinetic energy for the hydrodynamic case (left panel) with the spectrum for the dynamo case (right panel). We see that, in the hydrodynamic case, the contributions to high frequencies by the axial component (spectral index ) drop significantly faster than the contributions of the horizontal components and (spectral index ).
In the dynamo case, the spectral index of the horizontal components approaches which is close to the axial component, which practically does not change compared to the hydrodynamic case. This change is not particularly strong, but it is still clearly noticeable.
IV Conclusions
The kinematic models reveal the existence of two different branches for dynamo action. One branch is characterized by a highly localized and stationary dynamo that only occurs at rather large magnetic Reynolds number. The second type of dynamo is of oscillatory type with the magnetic field filling a larger proportion of the fluid domain. This type can occur at much lower magnetic Reynolds number, which, however requires suitable boundary conditions in terms of either a thin outer layer with large magnetic diffusivity or a comparable emulation with pseudo-vacuum BC. However, the kinematic model, which we assume best reflects the experimental setup, has a critical magnetic Reynolds number of , which is about twice as high as achievable in the planned precession dynamo experiment. Interestingly, this agrees quite well with the value derived by Goepfert and Tilgner [20] by means of simulations in a precessing cube.
We then applied direct numerical simulations and have calculated a comparative solution in order to investigate the time evolution of the velocity field and how it is influenced by the magnetic field through the Lorentz force and to check for consistentency with the kinematic models. The configuration of this DNS essentially corresponds to that of the kinematic model, in which a stationary outer layer with large magnetic diffusivity emulates the outer space in the laboratory. Unfortunately, it is not possible to achieve the large Reynolds numbers, as they are relevant in the experiment, due to the limitations of the numerics and the hardware, respectively. This is mainly due to the extremely thin boundary layers, which require a sufficiently high resolution. As a consequence, we find dynamo solutions in the DNS only for relatively large magnetic Reynolds numbers so that we effectively deal with a magnetic Prandtl number of the order of one with the magnetic field being small-scale and fluctuating on short time scales. At this stage it cannot be concluded with certainty that this behavior will change if we move to larger Reynolds numbers, as there are currently insufficient parameter studies available for such scaling. For the parameters considered, a small-scale burst dynamo is the typical solution. These bursts are characterized by peaks in magnetic energy, which, as previously mentioned, represent the sudden excitation of large-scale magnetic field structures. Since these structures are isolated and not sustained throughout the entire volume, they are rapidly destroyed by strong magnetic diffusion. This is followed by a period without the excitation of large-scale structures (no peaks in magnetic energy) until the hydromagnetic dynamo once again excites some large-scale magnetic field structures. Our results align with findings [36, 37, 38], which conclude that increasing the magnetic Prandtl number leads to a transition from large-scale dynamos to small-scale dynamos. Thus, at high values of , small-scale dynamos predominate, which is consistent with our case.
To summarize, the characteristics of the magnetic fields from the kinematic simulations are clearly different from those of the DNS. However, if the magnetic field from the DNS is averaged over a sufficiently long period of time so that the occasional occurrence of the strongly localized patches is also averaged out, the structure is comparable to the structure of the weakly efficient branch in the kinematic models, which only appears at large magnetic Reynolds numbers. This explains, at least in part, the high magnetic Reynolds number required to obtain a dynamo in the DNS.
Averaging the magnetic fields over a sufficient period of time, the small scale magnetic field structures give a similar pattern as for the weak branch of the kinematic solutions, while the occasional bursts appear to have a similar character to the strong branch, but fail to establish a corresponding dynamo solution.
Acknowledgements.
This work benefited from support through Project Nos. GR 967/7-1 and GI 1405/1-1 of the Deutsche Forschungsgesellschaft (DFG). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). Ján Šimkanin is grateful for support from the Alexander-von-Humboldt Stiftung (CZE 1079936).References
- Stieglitz and Müller [2001] R. Stieglitz and U. Müller, “Experimental demonstration of a homogeneous two-scale dynamo,” Phys. Fluids 13, 561–564 (2001).
- Gailitis et al. [2000] A. Gailitis, O. Lielausis, S. Dement’ev, E. Platacis, A. Cifersons, G. Gerbeth, T. Gundrum, F. Stefani, M. Christen, H. Hänel, and G. Will, “Detection of a Flow Induced Magnetic Field Eigenmode in the Riga Dynamo Facility,” Phys. Rev. Lett. 84, 4365–4368 (2000).
- Monchaux et al. [2007] R. Monchaux, M. Berhanu, M. Bourgoin, M. Moulin, P. Odier, J.-F. Pinton, R. Volk, S. Fauve, N. Mordant, F. Pétrélis, A. Chiffaudel, F. Daviaud, B. Dubrulle, C. Gasquet, L. Marié, and F. Ravelet, “Generation of a Magnetic Field by Dynamo Action in a Turbulent Flow of Liquid Sodium,” Phys. Rev. Lett. 98, 044502 (2007), 10.1103/PhysRevLett.98.044502.
- Giesecke et al. [2018] A. Giesecke, T. Vogt, T. Gundrum, and F. Stefani, “Nonlinear Large Scale Flow in a Precessing Cylinder and Its Ability To Drive Dynamo Action,” Phys. Rev. Lett. 120, 024502 (2018).
- Stefani et al. [2019] F. Stefani, A. Gailitis, G. Gerbeth, A. Giesecke, T. Gundrum, G. Rüdiger, M. Seilmayer, and T. Vogt, “The dresdyn project: liquid metal experiments on dynamo action and magnetorotational instability,” Geophys. Astrophys. Fluid Dyn. 113, 51 – 70 (2019).
- Léorat [2006] J. Léorat, “Large scales features of a flow driven by precession,” Magnetohydrodynamics 42, 143–151 (2006).
- Malkus [1968] W. V. R. Malkus, “Precession of the Earth as the Cause of Geomagnetism,” Science 160, 259–264 (1968).
- Cébron et al. [2019] D. Cébron, R. Laguerre, J. Noir, and N. Schaeffer, “Precessing spherical shells: flows, dissipation, dynamo and the lunar core,” Geophys. J. Int. 219, S34–S57 (2019).
- Olson [2013] P. Olson, “The new core paradox,” Science 431–432, 013008 (2013).
- Landeau et al. [2022] M. Landeau, A. Fournier, H.-C. Nataf, D. Cébron, and N. Schaeffer, “Sustaining earth’s magnetic dynamo,” Nature Reviews Earth and Environment 3, 255 – 269 (2022).
- Rochester et al. [1975] M. G. Rochester, J. A. Jacobs, D. E. Smylie, and K. F. Chong, “Can precession power the geomagnetic dynamo?” Geophys. Journal 43, 661–678 (1975).
- Kerswell [1996] R. R. Kerswell, “Upper bounds on the energy dissipation in turbulent precession,” J. Fluids Mech. 321, 335–370 (1996).
- Pizzi et al. [2021] F. Pizzi, A. Giesecke, J. Šimkanin, and F. Stefani, “Prograde and retrograde precession of a fluid -filled cylinder,” New J. Phys. 23, 123016 (2021).
[…]
- Giesecke et al. [2024] A. Giesecke, T. Vogt, F. Pizzi, V. Kumar, F. Garcia-Gonzalez, T. Gundrum, and F. Stefani, “The global flow state in a precessing cylinder.” J. Fluid Mech. XXX, in press (2024).
- Stys and Dumberry [2020] C. Stys and M. Dumberry, “A past Lunar dynamo thermally driven by the precession of its inner core,” Journal of Geophysical Research: Planets 125, e2020JE006396 (2020).
[…]
- Wu and Roberts [2009] C.-C. Wu and P. Roberts, “On a dynamo driven by topographic precession,” Geophys. Astrophys. Fluid Dyn. 103, 467–501 (2009).
- Nore et al. [2011] C. Nore, J. Léorat, J.-L. Guermond, and F. Luddens, “Nonlinear dynamo action in a precessing cylindrical container,” Phys. Rev. E 84, 016317 (2011).
- Goepfert and Tilgner [2016] O. Goepfert and A. Tilgner, “Dynamos in precessing cubes,” New J. Phys. 18, 103019 (2016), 10.1088/1367-2630/18/10/103019.
- Gans [1971] R. F. Gans, “On hydromagnetic precession in a cylinder,” J. Fluids Mech. 45, 111–130 (1971).
- Giesecke et al. [2019] A. Giesecke, T. Vogt, T. Gundrum, and F. Stefani, “Kinematic dynamo action of a precession-driven flow based on the results of water experiments and hydrodynamic simulations.” Geophys. Astrophys. Fluid Dyn. 113, 235–255 (2019).
- Pizzi, Giesecke, and Stefani [2021] F. Pizzi, A. Giesecke, and F. Stefani, “Ekman boundary layers in a fluid filled precessing cylinder,” AIP Advances 11, 035023 (2021).
- Blackburn and Sherwin [2004] H. M. Blackburn and S. J. Sherwin, “Formulation of a Galerkin spectral element-Fourier method for three-dimensional incompressible flows in cylindrical geometries,” J. Comp. Phys. 197, 759–778 (2004).
- Blackburn et al. [2019] H. M. Blackburn, D. Lee, T. Albrecht, and J. Singh, “Semtex: a spectral element–Fourier solver for the incompressible Navier–Stokes equations in cylindrical or Cartesian coordinates,” Comput. Phys. Cmm. 245, 106804–1–13 (2019).
- Giesecke, Stefani, and Gerbeth [2008] A. Giesecke, F. Stefani, and G. Gerbeth, “Kinematic simulation of dynamo action by a hybrid boundary-element/finite-volume method,” Magnetohydrodynamics , 237 – 252 (2008).
- Giesecke et al. [2010] A. Giesecke, C. Nore, F. Stefani, G. Gerbeth, J. Léorat, F. Luddens, and J.-L. Guermondz, “Electromagnetic induction in non-uniform domains,” Geophys. Astrophys. Fluid Dyn. 104, 505 – 529 (2010).
- Giesecke [2024a] A. Giesecke, “supplemental material,” www.giesecke.net (2024a).
- Giesecke [2024b] A. Giesecke, “supplemental material,” www.giesecke.net (2024b).
- Rüdiger, Elstner, and Ossendrijver [2003] G. Rüdiger, D. Elstner, and M. Ossendrijver, “Do spherical -dynamos oscillate?” Astron. Astrophys. 406, 15–21 (2003).
- Wilbert [2024] M. Wilbert, Implementation and application of a pseudo-spectral MHD solver combined with an immersed boundary method to support the DRESDYN dynamo experiment, Ph.D. thesis, Fakultät für Physik und Astronomie der Ruhr-Universität Bochum (2024).
- Kreuzahler et al. [2014] S. Kreuzahler, D. Schulz, H. Homann, Y. Ponty, and R. Grauer, “Numerical study of impeller-driven von Kármán flows via a volume penalization method,” New J. Phys. 16, 103001 (2014), 10.1088/1367-2630/16/10/103001.
- Kreuzahler et al. [2017] S. Kreuzahler, Y. Ponty, N. Plihon, H. Homann, and R. Grauer, “Dynamo enhancement and mode selection triggered by high magnetic permeability,” Phys. Rev. Lett. 119, 234501 (2017).
- Miralles, Plihon, and Pinton [2015] S. Miralles, N. Plihon, and J.-F. Pinton, “Lorentz force effects in the bullard–von kármán dynamo: saturation, energy balance and subcriticality,” J. Fluids Mech. 775, 501–523 (2015).
- Subramanian [2003] K. Subramanian, “Hyperdiffusion in nonlinear large- and small-scale turbulent dynamos,” Phys. Rev. Lett. 90, 245003 (2003).
- Etchevest, Fontana, and Dmitruk [2022] M. Etchevest, M. Fontana, and P. Dmitruk, “Behavior of hydrodynamic and magnetohydrodynamic turbulence in a rotating sphere with precession and dynamo action,” Physical Review Fluids 7, 103801 (2022).
- Richardson, Hollerbach, and Proctor [2012] K. Richardson, R. Hollerbach, and M. Proctor, “From large-scale to small-scale dynamos in a spherical shell,” Phys. Fluids 24, 107103 (2012).
- Hughes, Cattaneo, and Thelen [2001] D. Hughes, F. Cattaneo, and J.-C. Thelen, Dynamo and Dynamics, a Mathematical Challenge, NATO Science Series (Kluwer Academic Publishers, Dordrecht / Boston / London, 2001).

### v3

@@
 We compare a simplified kinematic approach based on the solution of the magnetic induction equation with a prescribed velocity field with the results from a self-consistent three-dimensional simulation of the complete set of magnetohydrodynamic equations.
-The different kinematic models show a strong influence of the electromagnetic properties of outer layers, which manifests itself in the existence of two different branches with dynamo action.
 In all cases, we observe a minimum for the onset of dynamo action in a transitional regime, within which the hydrodynamic flow undergoes a change from a large-scale to a more small-scale, turbulent behaviour.
-However, significant differences in the absolute values for the critical magnetic Reynolds number occur, and they can be explained by the low rotation rate in the MHD simulations achievable for technical reasons.
-In these models the character of the dynamo solution is small-scale and the magnetic energy remains significantly smaller than the kinetic energy of the flow.
+However, significant differences in the absolute values for the critical magnetic Reynolds number occur depending on the physical properties of the external layers surrounding the flow active domain.
+The strong influence of the electromagnetic properties of outer layers with the large variation of the critical magnetic Reynolds number can be related to the existence of two different branches with dynamo action.
+In contrast to the kinematic models, the nonlinear MHD simulations reveal a small scale dynamo solution with the magnetic energy remaining significantly smaller than the kinetic energy of the flow.
 In irregular intervals, we observe dynamo bursts with a local concentration of the magnetic field, resulting in a global increase of the magnetic energy by a factor of 3 to 5.
-However, diffusion of the local patches caused by strong local shear is too rapid, causing these features to exist for only a short period (i.e., shorter than a rotation period) so that their dynamical impact on the dynamo remains small.
+However, diffusion of the local patches caused by strong local shear is too rapid, causing these features to exist for only a short period so that their dynamical impact on the dynamo remains small.
 As the magnetic field is small-scale and weak, the nonlinear feedback on the flow through the Lorentz force remains small and arises essentially in terms of a slight damping of the fast timescales, whereas there is no noticeable change in flow amplitude compared to the hydrodynamic case.
+A connection with the kinematic models can be derived by looking at the time-averaged field of the MHD dynamo solution.
+This is comparable to the eigenmode of the inefficient branch of the kinematic models, which explains their large critical magnetic Reynolds number.
 I Introduction
-The experimental investigation of magnetic field generation through a flow of electrically conductive fluid is of great interest, as such a dynamo process is crucial for many astrophysical bodies and allows conclusions to be drawn about their evolution and internal structure.
-Following the successful first-generation dynamo experiments in Karlsruhe [1] and Riga [2], which confirmed the fundamental principle, and subsequent dynamo experiments in Cadarache [3], characterized by a wide variety of different dynamic behaviours, a new experiment is currently being set up at HZDR [4, 5].
-In this experiment, a flow of liquid sodium in a precessing cylinder is intended to generate a magnetic field.
-A mechanical drive such as precession is indeed capable of generating strong flows, allowing for the large flow amplitudes required for the occurrence of the dynamo effect.
-This contrasts with potentially more straightforward mechanisms, such as convection, which at typical laboratory scales may not achieve sufficiently vigorous flows due to lower driving efficiency.
-In addition, precession as a flow drive makes perfect sense in a dynamo experiment because driving on the largest possible length scale – the size of the container – opens up significantly more degrees of freedom for establishing a (turbulent) flow through cascade processes than in the case of the previously highly optimized experiments[6].
-Precession is a repeatedly proposed driving mechanism for natural dynamos, whether for Earth [7] or the Moon [8].
-Paleomagnetic records suggest that the geodynamo was active even before the formation of the Earth’s solid inner core, which implies that thermochemical convection, driven by the cooling and crystallization of the core, was not always the primary mechanism.
-Before the onset of thermochemical convection, alternative mechanisms must have driven the early geodynamo.
-Among the most plausible candidates are precession-driven and tide-driven flows, which could have provided the necessary energy for maintaining the magnetic field in the Earth’s early history.
-This follows from recent experiments and numerical computation yield rather large values for the heat conductivity under conditions typical for the Earth’s liquid core, which imposes strict energetic constraints for the convective state, particularly for the early geodynamo (before the formation of the solid inner Earth’s core, see e.g.
+The experimental investigation of magnetic field generation through a flow of electrically conductive fluid is of great interest, as such a magnetohydrodynamic dynamo process is crucial for many astrophysical bodies and allows conclusions about their evolution and internal structure.
+Following the successful first-generation dynamo experiments in Karlsruhe [1] and Riga [2], which confirmed the fundamental principle, subsequent dynamo experiments in Cadarache (Monchaux et al.
+[3]) revealed a wide variety of different dynamic behaviours.
+Currently, a new experiment is being set up at HZDR (Stefani et al.
+[4]).
+In the DRESDYN experiment, a flow of liquid sodium in a precessing cylinder is intended to generate a magnetic field.
+Precession is a repeatedly proposed driving mechanism for natural dynamos, whether for the Moon [5] or for the Earth [6].
+Regarding the Lunar dynamo, which around 4 billion years ago generated a magnetic field of comparable strength to that of today’s Earth, precession is indeed the most promising candidate to explain the observations including the disappearance of the dynamo about 3 billion years ago (see e.g.
+Stys and Dumberry [7]).
+One of the main reasons to consider precession driven dynamo action for the geodynamo are paleomagnetic records, which suggest that the geodynamo was active even before the formation of the Earth’s solid inner core.
+This, in turn, implies that thermochemical convection, driven by the cooling and crystallization of the core, was not always the primary mechanism and before the onset of thermochemical convection, alternative mechanisms must have driven the early geodynamo.
+Support for this hypothesis comes from recent experiments and numerical computations that yield rather large values for the heat conductivity under conditions typical for the Earth’s liquid core, which imposes strict energetic constraints for the convective state, particularly for the early geodynamo before the formation of the solid inner Earth’s core (see e.g.
 Olson [8], Landeau et al.
 [9]).
-While it is relatively clear that the laminar flow directly driven by precession is energetically insignificant [11], the original idea of Malkus was based on a turbulent base state [7], which was underpinned by estimates of an upper limit to (viscous) dissipation by Kerswell [12].
-Indeed, precession-driven fluid flows exhibit several instabilities, allowing both forward and backward cascades of energy transfer [13] that end up in a flow state with small-scale turbulence superimposed by large-scale flow contributions essentially in terms of a circulation opposite to the original rotation of the container [14, 15].
-Regarding the Lunar dynamo, which around 4 billion years ago generated a magnetic field of comparable strength to that of today’s Earth, precession is indeed the most promising candidate to explain the observations including the disappearance of the dynamo about 3 billion years ago, see e.g.
-Stys and Dumberry [16].
-Precession-driven dynamos have been the subject of numerous numerical studies, and there are now successful dynamo simulations in all conceivable geometries, including spheres [17], spheroids [18], cylinders [19], or cubes [20].
-A preliminary experiment with sodium in a precessing cylinder was conducted in the early 1970s [21].
-This experiment demonstrated the amplification of an externally applied magnetic field by a factor of three, however, the setup at that time was not large enough to achieve the magnetic Reynolds numbers required for a dynamo.
-Preliminary investigations using a kinematic model based on the geometry of the planned experiment and the large-scale flow from hydrodynamic simulations showed that the dynamo effect works best in the narrow transition region that separates the subcritical from the supercritical state [4, 22].
-In the present study, we deepen the investigation of the dynamo effect in the transitional regime and furthermore, we investigate the influence of the magnetic boundary conditions and turn to full self-consistent models to study the influence of the nonlinear terms and the Lorentz force feedback on the flow.
+While it is relatively clear that the laminar flow directly driven by precession is energetically insignificant (Rochester et al.
+[10]), the original idea of Malkus was based on a turbulent base state[6], which was later underpinned by estimates of an upper limit to (viscous) dissipation by Kerswell [11].
+Indeed, precession-driven fluid flows exhibit several instabilities, allowing both forward and backward cascades of energy transfer (Pizzi et al.
+[12]) that end up in a flow state with small-scale turbulence superimposed by large-scale flow contributions essentially in terms of a circulation opposite to the original rotation of the container (Wilbert, Giesecke, and Grauer [13], Giesecke et al.
+[14]).
+This behaviour has also been observed in the laboratory[15] and so far various experiments have shown that precession is capable of generating strong flows, principally allowing for the large flow amplitudes required for the occurrence of the dynamo effect[16, 17, 18, 19, 20, 21, 22, 23, 24].
+This contrasts with potentially more straightforward mechanisms, such as convection, which at typical laboratory scales may not achieve sufficiently vigorous flows due to lower driving efficiency [25, 26, 27].
+Precession-driven dynamos also have been the subject of numerous numerical studies, and there are now successful dynamo simulations in all conceivable geometries, including spheres [28], spheroids [29], or cubes [30].
+Nore et al.
+[31] and Cappanera et al.
+[32] investigated dynamo solutions in a cylindrical geometry comparable to the DRESDYN experiment, but with small Reynolds numbers () and a rather large precession ratio of .
+The induced magnetic fields showed a complex structure, which correspond most closely to a quadrupolar geometry.
+Preliminary investigations with particular regard to the DRESDYN dynamo experiment utilized a kinematic model based on the geometry of the planned experiment and the large-scale flow from hydrodynamic simulations (Giesecke et al.
+[33, 34]).
+These models showed that the dynamo effect works best in the narrow transition region that separates the subcritical from the supercritical state (Pizzi et al.
+[35], Kumar et al.
+[36]).
+In the present study, we deepen the investigation of the dynamo effect in the transitional regime, investigate the influence of outer non-fluid layers, and compare kinematic models with a self-consistent magnetohydrodynamic (MHD) model to study the influence of the nonlinear terms and the Lorentz force feedback on the flow.
 The present study is organized as follows.
-In section II.1, we introduce the equations that describe an electrically conducting fluid flow in a precessing cylinder and the related magnetic induction.
-We further summarize the essential response of the fluid to the forcing imposed by the precession of the container and present the evolution when increasing this forcing.
-The subsequent section II.3 is dedicated to the results obtained from kinematic dynamo models, where only the induction equation is solved numerically assuming a prescribed, time-independent flow.
+In section II, we introduce the equations that describe an electrically conducting fluid flow in a precessing cylinder and the related magnetic induction.
+We further summarize the essential response of the fluid to the forcing imposed by the precession of the container and present the evolution when the forcing is increased.
+The subsequent section III is dedicated to the results obtained from kinematic dynamo models, where only the induction equation is solved numerically assuming a prescribed, time-independent flow.
 In this section, we detail the impact of the electrical conductivity of the container.
-Section III focuses on the corresponding results of the full set of magnetohydrodynamic (MHD) equations and reveals differences and similarities with the kinematic dynamo.
-We end our study with the conclusions in section IV, where we try to adopt our results for the forthcoming dynamo experiment at HZDR.
-II Kinematic approach
-II.1 Setup & Equations
-Initially, we focus on the kinematic problem, where only the magnetic induction equation is solved and a time-independent velocity field is prescribed, which is supposed to represent the mean (time-averaged) velocity in the experiment.
-The main finding from this model is a value for the critical magnetic Reynolds number, which is related to the minimum amplitude of the velocity field required for a dynamo to occur.
-The kinematic approach does neither consider the back-reaction of the magnetic field on the fluid flow via the Lorentz force so that we have no possibility for any saturation, nor the impact of small-scale fluctuations that could hamper (turbulent diffusion) or support (-effect) the dynamo process.
+Section IV focuses on the corresponding results obtained from the full set of magnetohydrodynamic (MHD) equations and reveals differences and similarities with the kinematic dynamo.
+We end our study with the conclusions in section V, where we try to adopt our results for the forthcoming dynamo experiment at HZDR.
+II Setup of the kinematic dynamo models
+II.1 Equations and numerical methods
+Initially, we focus on the kinematic problem, where only the magnetic induction equation is solved, and a time-independent velocity is prescribed, which is supposed to represent the mean (time-averaged) velocity in the experiment.
+Here we compare different configurations without and with outer layers of a conductive material surrounding the actual medium on which the force is exerted by the precession.
+The main finding from these models is a value for the critical magnetic Reynolds number, which is related to the minimum amplitude of the velocity field required for a dynamo to occur.
+The kinematic approach does neither consider the back-reaction of the magnetic field on the fluid flow via the Lorentz force nor the impact of small-scale fluctuations that could obstruct (turbulent diffusion) or support (-effect) the dynamo process.
 We calculate the temporal evolution of the magnetic flux density (in the following called ’magnetic field’) by numerically solving the magnetic induction equation, which reads
 In Eq.
-(1) represents the velocity field and is the magnetic Reynolds number defined with the angular velocity of the cylinder wit radius and the magnetic diffusivity , which is related to the electrical conductivity by (see also Figure 1a).
-In all cases, the inner region of the simulation domain consists of a cylindrical container with radius and height (in dimensionless unities) which is filled with an electrically conducting fluid with a magnetic diffusivity .
-The fluid flow is prescribed by the data taken from hydrodynamic simulations presented in Pizzi, Giesecke, and Stefani [23], Pizzi et al.
-[13], Giesecke et al.
-[15].
-In these studies, it is found that the flow in a cylinder emerging at a large nutation angle can be represented by the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow.
-A more detailed prescription of the structure of the flow field is given in the following subsection.
-The flow fields used in this study are derived from hydrodynamic simulations with the spectral element code SEMTEX[24, 25].
-A detailed description of the flow field as a function of the precession ratio defined by can be found in [23, 13, 14, 15].
-In the following, we briefly summarize the main properties that are relevant for our present study.
-The flow in a cylinder precessing at a large nutation angle can be characterized by three regimes.
-In all cases, it is possible to represent the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow.
-The structure of the time-averaged flow fields as applied in the kinematic model is shown in Figure 2 for three precession ratios that are characteristic for the three flow regimes.
-In the subcritical regime, the flow is essentially laminar, and the geometry can be described by the superposition of the rotational motion due to the rotating cylinder and the directly driven flow (essentially an mode proportional to , see left column in Figure 2).
-Overlaying this is the zonal flow maintained by the cylinder’s rotation (geostrophic and axisymmetric).
-Without precession, this would be a pure solid-body rotation with with but due to the influence of precession the radial profile of the zonal flow undergoes a significant modification which leads to a significant deceleration of the fluid rotation in the bulk (see bottom row in Figure 2).
-For larger forcing the flow transients into a supercritical region, in which the large-scale flow in the inner region has almost completely stopped so that the fluid essentially performs small-scale turbulent fluctuations and large-scale components are then only found near the walls of the container (see right column in Figure 2).
-These two regimes are linked by a transition zone, which is characterized by an intermittent behaviour of the large-scale flow with elements from both regimes.
-In this transitional region, there exists an additional large-scale contribution in the form of a double roll pattern as shown in the central column of Figure 2 (third row).
-II.2 Code
-II.2.1 Algorithm
+(1) represents the (incompressible) velocity field and is the magnetic Reynolds number defined with the angular velocity of the cylinder, the radius and the magnetic diffusivity (see also Figure 1a).
 Assuming a prescribed time-independent flow field, Eq.
@@
 The scheme applies the constraint transport method to guarantee divergence-free solutions and is described in more detail in Giesecke, Stefani, and Gerbeth [37].
-Extensive tests of the algorithm, in particular the impact of locally varying material properties, such as electrical conductivity and comparison with an alternative approach based on Spectral/Finite Elements (the SFEMaNS code) can be found in Giesecke et al.
+Extensive tests of the algorithm, in particular the impact of locally varying material properties, such as electrical conductivity and comparison with an alternative approach based on Spectral/Finite Elements (the SFEMaNS code[38]) can be found in Giesecke et al.
 [39].
-II.2.2 Boundary conditions
-We apply pseudo-vacuum conditions for the magnetic field, which read (, where is the unit normal vector at the boundary so that the tangential components of the magnetic field an the outer boundary vanish.
-It is well known that pseudo-vacuum conditions in kinematic models usually underestimate the critical magnetic Reynolds number that must be exceeded for the dynamo to start.
-The reduction depends on the geometry, i.e.
-essentially on the aspect ratio, and amounted, for example, to up to 30% in the kinematic models of the VKS dynamo compared to insulator boundary conditions [27].
-Nevertheless, we must restrict ourselves to the application of pseudo-vacuum conditions at the outer boundary of our system, because the strong shear near the outer boundaries enforces a high spatial resolution, which rules out the application of the Boundary-Element-Method (BEM, see e.g.
-Giesecke, Stefani, and Gerbeth [26]) for the consideration of insulating boundary conditions, as this method requires a large fully populated matrix to be kept in the memory.
-The size of this matrix scales with , where represents the number of grid points on the cylinder surface so that even current computers with a memory capacity of more than 1 TB are not sufficient to calculate the corresponding models.
-Instead, we estimate the impact of the magnetic boundary conditions employing various setups with one or two outer layers with thickness and as sketched in Figure 1.
-The underlying idea is that a sufficiently thick boundary layer with poor electrical conductivity (large magnetic diffusivity) provides a good shielding of the inner volume filled with the electrically conductive fluid so that the actual nature of the electrical boundary conditions sufficiently far away from the induction region is no longer important.
+II.2 Boundary conditions and nested outer layers
+We estimate the impact of the magnetic boundary conditions employing various setups with one or two outer layers with thickness and as sketched in Figure 1b.
+The underlying idea is that a sufficiently thick outer layer with poor electrical conductivity (large magnetic diffusivity) reduces the impact of the magnetic boundary conditions on the inner volume filled with the electrically conductive fluid.
+We consider several models that differ in the presence of up to two outer layers with thicknesses and that enclose the internal region of the container filled with an electrically conducting fluid (Figure 1b).
+In all cases, the inner region of the simulation domain consists of a cylindrical container with radius and height (in dimensionless units) which is filled with an electrically conducting fluid with a magnetic diffusivity .
 We examine three different geometric configurations.
@@
 ) and two different values for the thickness of the wall layer.
-In the thin layer model we use with a resolution of grid cells and in the thick layer model we use with a resolution of grid cells, whereby the magnetic diffusivity in the outer layer remains fixed at the maximum value of .
-Finally, in the fourth model we again apply a thin cylinder wall with and add a second outer (non-rotating) layer with and , which is supposed to resemble a non-conducting exteriour.
-II.3 Results of the kinematic models
-II.3.1 General overview
-Previous kinematic models demonstrated that dynamo action is most effective when utilizing flow from the transitional region between the subcritical and supercritical states.
-A detailed investigation of which components of this realization are of decisive importance for the dynamo effect was carried out in Giesecke et al.
-[22] and showed that shear adjacent to the sidewall plays an important role as well as the presence of the double roll mode (DRM) which seems to be responsible for the considerable drop in the critical magnetic Reynolds number.
-II.3.2 No outer layers
+In the thin layer model we use with a resolution of grid cells and in the thick layer model we use with a resolution of grid cells, whereby the magnetic diffusivity in the wall layer remains fixed at the maximum value of .
+As we tackle the problem in the precession frame of reference (with the observer co-rotating with the precession while looking at the spinning container), the wall layer follows a rotational motion with and .
+Finally, in the fourth model we again apply a thin wall layer with and surrounded by a second, outer layer with and , which is supposed to resemble a non-conducting exterior (see Rädler et al.
+[40] and present discussion on page 14).
+This layer does not co-rotate with the container so that the velocity is set to zero within this layer.
+At the outer boundaries of the computational domain we apply pseudo-vacuum conditions for the magnetic field, which read (, where is the unit normal vector at the boundary so that the tangential components of the magnetic field an the outer boundary vanish.
+It is well known that pseudo-vacuum conditions in kinematic models usually underestimate the critical magnetic Reynolds number that must be exceeded for the dynamo to start.
+The reduction depends on the geometry, i.e.
+essentially on the aspect ratio, and amounted, for example, to up to 30% in the kinematic models of the VKS dynamo compared to realistic insulator boundary conditions (Giesecke et al.
+[39]).
+II.3 Velocity field
+The fluid flow is prescribed using the data taken from hydrodynamic simulations presented in Pizzi, Giesecke, and Stefani [41], Pizzi et al.
+[12], Wilbert, Giesecke, and Grauer [13] and Giesecke et al.
+[14].
+In these studies, it was found that the flow in a cylinder emerging at a large nutation angle can be represented by the time-averaged flow with a few large-scale inertial waves that capture the major part of the kinetic energy of the flow.
+In dependence of the precession ratio defined by the flow basically can be characterized by three regimes, which we simply name the subcritical and the supercritical state, and the transition section between both states.
+A compact representation of the three regimes, which illustrates how the transition between the subcritical state and the supercritical state takes place, is provided in Figure 2, which shows the ratio of the energy of the turbulent flow component in the form of the time- and volume-averaged root-mean-square velocity to the energy of the actual axisymmetric geostrophic component as a function of .
+The plot illustrates the abrupt shift from the rotation-dominated regime (below ) to the turbulence-dominated regime (above ).
+However, although the fluctuations of the flow with respect to the fluid rotation increase strongly with , the dominant energetic contributions of the precession-induced flow can always be captured by a few large-scale inertial modes, since the significant jump in Figure 2 is determined by both, the increase in turbulence intensity, and the deceleration of the fluid rotation in the core of the cylinder (see Giesecke et al.
+[14]).
+The typical structure of the time-averaged flow fields is shown in Figure 3 for three precession ratios that are characteristic for the three flow regimes.
+In the subcritical state (left column in Figure 3), the flow is essentially laminar, and the geometry can be described by the directly driven flow (first and second row, essentially an mode proportional to ) and a zonal flow maintained by the cylinder’s rotation (geostrophic and axisymmetric, third row).
+Without precession, this would be a pure solid-body rotation with but due to the influence of precession the radial profile of the zonal flow is significantly modified (i.e.
+deceleration of the fluid rotation in the bulk, see bottom row in Figure 3).
+For larger forcing the flow transients into a supercritical region, in which the large-scale flow in the inner region almost vanishes so that the fluid essentially performs small-scale turbulent fluctuations and large-scale components are only found near the walls of the container (see right column in Figure 3).
+These two regimes are linked by a transition zone, which is characterized by an intermittent behavior of the large-scale flow with elements from both regimes.
+In this transitional region, there exists an additional large-scale contribution in the form of a double roll pattern as shown in the third row of the central column of Figure 3.
+This double roll is rather similar to the (time-averaged) poloidal velocity field in the VKS dynamo111In that dynamo experiment the flow was driven by two oppositely rotating disks, and it is well known that such a flow structure generates a dynamo with a relatively low critical magnetic Reynolds number[42] (when also the associated azimuthal flow is taken into account).
+Indeed, previous kinematic models using a time-averaged flow driven by precession demonstrated that dynamo action is most effective when utilizing the flow obtained in the transitional region between the subcritical and supercritical states [33, 34].
+A detailed investigation of which components (in terms of inertial modes) are of decisive importance for the dynamo effect was carried out in Giesecke et al.
+[34] and showed that shear adjacent to the side walls plays an important role as well as the presence of the double-roll mode which seems to be responsible for the considerable drop in the critical magnetic Reynolds number.
+III Kinematic growth rates and critical magnetic Reynolds number
+III.1 General overview
+So far kinematic models for the DRESDYN dynamo experiment mostly applied pseudo-vacuum conditions for the magnetic field and only few models have been published where an outer wall with finite thickness and with different magnetic diffusivity has been considered [34].
+These models indicate that a container wall made of stainless steel, i.e.
+with a diffusivity that is larger by one order of magnitude when compared to the liquid sodium in the interior, yields growth rates for the magnetic energy similar to the models with pseudo-vacuum conditions.
+In the following, we will review and deepen these results.
+III.2 No layer
 The typical behavior of the temporal evolution of the magnetic energy is shown in Figure 4a.
-Here we focus on two paradigmatic cases with the flow obtained at scaled to a magnetic Reynolds number of (red curve) and (blue curve), which is slightly below and above the dynamo threshold.
-A striking feature is the oscillating behavior of the energy, which is superimposed on the exponentially growing part.
-The oscillating character of the solution can be clearly seen in the animation, which can be found in the supplementary material to this study [28].
-A systematic study of the impact of the velocity fields obtained at different the critical magnetic Reynolds number required for the onset of dynamo action which is shown in Figure 3b.
-We do not find any dynamos (at least up to ) for the flow fields in the laminar regime (i.e.
-below ).
-The lowest possible magnetic Reynolds number, is , which is obtained for the flow field from hydrodynamic simulations at .
-It is remarkable, that we find a significant smaller value for in the range than above, which coincides exactly with the transition regime of the velocity field from the subcritical to the supercritical state.
-We therefore assume that the axially symmetric double roll occurring exclusively in this parameter range plays a significant role for the onset of dynamo action at comparatively small magnetic Reynolds numbers.
-II.3.3 Thin outer layer
-In the following we focus on the velocity field that provides the minimum critical magnetic Reynolds number in our first models so that all results discussed for the kinematic approach in the reminder of this study are calculated from the velocity field obtained at (central column in Figure 2).
-We continue with the setup that considers an additional outer layer with thickness with increased magnetic diffusivity that emulates the container wall.
-If we look at the temporal development of the magnetic energy and its spatial distribution, we recognize two different types of dynamo action.
-If the wall diffusivity is close to that of the fluid (i.e.
+Here we focus on two paradigmatic cases with the flow obtained at scaled to a magnetic Reynolds number of (red curve, decaying solution) and (blue curve, growing solution), which are slightly below and above the dynamo threshold.
+A striking feature is the oscillating component of the energy, which is superimposed on the exponentially growing part.
+For a flow field with a slightly increased , the behavior looks similar in principle, but then a higher magnetic Reynolds number is required for the onset of a dynamo, and the frequency of the oscillating component increases as shown by the orange and green curves in Figure 4a, which result from the flow field obtained at .
+The distribution of the magnetic energy looks similar in both cases (see Figure 4b).
+It is worth mentioning in relation to results including outer boundary layers (see below) that a significant part of the magnetic field energy is observed in the bulk of the cylinder.
+The temporal dynamics due to the oscillating field component are shown in an animation, which can be found in the supplementary material to this study [43].
+III.3 Wall layer
+We continue with the setup that considers an additional wall layer with thickness .
+Within this layer the magnetic diffusivity is increased in order to emulate the properties of a container wall made of stainless steel.
+With such an outer layer, we essentially find two new features: (1) the onset for dynamo action is shifted to larger magnetic Reynolds numbers
+and (2) we can distinguish two different types of dynamo action: if the wall diffusivity is close to that of the fluid (i.e.
 ), a dynamo only occurs at very large ().
-This dynamo is time-independent (no oscillatory components, see orange and green curve in Figure 4a) and the associated field energy is essentially concentrated in a narrow region near the end caps (see iso-surfaces in Figure 4c).
-If the outer layer has an larger than (see blue curve in Figure 4b), we obtain an oscillating solution that already emerges at much smaller , which is exemplified by the blue curve for in Figure 4a.
-In that case the magnetic field energy is widely distributed throughout the entire volume (however still with a characteristic pattern, see Figure 4d.) The impact of the outer layer is illustrated by the growth rates shown in Figure 4b.
-When the magnetic diffusivity of the outer layer is increased to emulate the effect of a stainless steel container, an eigenmode is excited, the growth rates of which are represented by the inverted parabola in the individual curves in Figure 4b.
-This special shape results in positive growth rates for and for further increasing diffusivities the behaviour increasingly approaches that of pseudo-vaccum boundary conditions without outer layer (see black curve).
-It is therefore the eigenmode that belongs to the inverse parabolic growth rate, which corresponds to a dynamo that occurs at relatively low achievable in the experiment, and causes a magnetic field that permeates the entire volume.
-Interestingly, this eigenmode belongs to an oscillating dynamo while the much less efficient dynamo is stationary, although the opposite behavior is generally assumed [30].
-II.3.4 Thick outer layer
-The beneficial effect of the outer layer with sufficiently large diffusivity vanishes, when its thickness is increased, as can be recognized in Figure 5a which shows the evolution of the magnetic energy for while the magnetic diffusivity remains at .
-Here we end up with a critical magnetic Reynolds number slightly above and the solution is again characterized by a non-oscillatory behaviour with the magnetic energy constraint to two small section close to the endcaps, which clearly indicates that we are in the low efficiency branch (see Figure 5c).
-Obviously, the dominant solution in this case is again given by the less efficient branch.
-Nevertheless, the highly efficient branch can be achieved again by assuming a more realistic configuration, where the container wall is emulated by a thin co-rotating layer with , while the laboratory is modeled by a thick outer layer with large diffusivity (actually, this should be an insulator in order to emulate air).
-In this case, we obtain an oscillating solution that occupies the entire cylinder volume and has a similar structure as in the case of the thin layer or pseudo-vacuum boundary conditions without an outer layer (see Figures 5b and 5d).
-However, the critical magnetic Reynolds number of is then about two and a half times as large as in these cases.
-Thus we see from the kinematic linear models that there are two different branches of dynamo solutions, the realization of which depends on the configuration of the outer boundary regions, i.e.
-essentially on the boundary conditions.
-We have not investigated here at which exact parameters the transition from one branch to the other takes place, since the experimental realization must in any case start from a basically infinitely large outer volume of an insulator, namely the air in the laboratory.
-Interestingly, the resulting critical magnetic Reynolds number of is very close to the value derived by Goepfert and Tilgner [20] for the precession dynamo with the help of simulations in a cube.
-In the following, we now want to solve the induction equation coupled to the Navier-Stokes equation with a comparable setup in a self-consistent MHD model, so that the time evolution of magnetic field and velocity field is solved simultaneously, including the feedback via the Lorentz force.
-III Full set of magnetohydrodynamic equations
-III.1 Code scheme
-An extension of the kinematic problem arises when we additionally consider the temporal evolution of the flow field.
-In this case, the Navier-Stokes equation needs to be solved as well, which in the case of precession-driven flow takes the form
-In addition to the terms commonly encountered in regular flow problems describing nonlinear interactions and dissipation, we have three contributions: the Coriolis force due to rotation, the Lorentz force due to the interaction with the magnetic field, and the Poincaré force accounting for acceleration due to perpetual change of the orientation of the cylinder’s axis of rotation.
+This dynamo has no oscillatory components (see orange and green curve in Figure 5a) and the associated field energy is concentrated in a narrow region near the end caps (top plot with iso-surfaces in Figure 5b).
+If the wall layer has an (see red and blue curves in Figure 5a), we obtain an oscillating solution that is similar to the solution in the previous case with pseudo-vacuum boundary conditions.
+In that case the magnetic field energy is again widely distributed throughout the entire volume with the characteristic pattern shown in the bottom plot of Figure 5b.
+A detailed view on the impact of the diffusivity of the outer layer on the growthrates is shown in Figure 6.
+When the magnetic diffusivity of the outer layer is increased to emulate the effect of a stainless steel container, an eigenmode is excited, the growth rates of which are represented by the inverted parabola in the individual colored curves in Figure 6.
+For this special shape results in positive growth rates at far below the point where the "regular" curve becomes positive.
+For further increasing diffusivities the behaviour approaches that of pseudo-vacuum boundary conditions without outer layer (illustrated by the black curve).
+It is therefore the eigenmode that belongs to the inverse parabolic growth rate, which corresponds to a dynamo that occurs at relatively low and causes a magnetic field that permeates the entire volume.
+The beneficial effect of the wall layer with (sufficiently) large vanishes, when its thickness is increased, as shown in Figure 7a which presents the evolution of the magnetic energy for while the magnetic diffusivity remains at .
+Here we end up with a critical magnetic Reynolds number slightly above and the solution is again characterized by a non-oscillatory behavior with the magnetic energy constraint to two small section close to the endcaps (see Figure 7b), indicating that we are in the low efficiency branch.
+III.4 Wall layer plus outer layer
+The high-efficient branch can also be achieved by assuming a configuration, where the container wall is emulated by a thin co-rotating layer with , while the exterior domain is modeled by a thick resting outer layer with .
+Here both additional layers have the same diffusivity, i.e.
+(actually, since the thick outer layer emulates the exterior this should be an insulator with ).
+In this case, we again obtain an oscillating solution that occupies the entire cylinder volume (see Figure 8) and has a similar structure as in the case of the thin layer or pseudo-vacuum boundary conditions without an outer layer shown in Figures 4 and 5.
+However, the critical magnetic Reynolds number of is now about two and a half times as large as in the previous cases and the amplitude of the oscillating part remains rather small.
+III.5 Impact of the structure of the velocity field
+In the following we focus on the setup with wall layer and outer layer for which we applied various time averaged velocity fields obtained within the interval .
+It is known from the hydrodynamic investigations that the structure of the velocity field changes only slightly within the transition between subcritical and supercritical state.
+Nevertheless, these gradual changes are sufficient to cause a variation in the critical magnetic Reynolds number, especially at the borders of the transition region.
+Figure 9 shows the growth rates for the model with wall and outer layers but varying the flow structure by using time-averaged flow fields obtained at different .
+Similar to the results with wall layer only (Figure 6), we see a behavior for some velocity fields that can be described by an inverse parabola, at least around the range with positive growth rates.
+This does not apply to velocity fields with or with .
+For these parameters, the parabolic behavior does not occur, and positive growth rates are found - if at all - for very large .
+The intersection of the growth rates with the abscissa reveals the critical magnetic Reynolds number required for the onset of a dynamo.
+Figure 10 shows this number as a function of for two different setups.
+The blue curve shows for the original setup without any outer layer.
+The most outstanding feature of the blue curve is the reduction of the critical magnetic Reynolds number in the range with the lowest value being , obtained for the flow field from hydrodynamic simulations at .
+Below no more dynamos were found, while above the critical magnetic Reynolds number is in the range of about .
+The region coincides exactly with the transition region from the subcritical to the supercritical state of the velocity field, and we therefore assume that the axially symmetric double roll occurring exclusively in this parameter range plays a significant role for the onset of dynamo action at comparatively small magnetic Reynolds numbers (see Giesecke et al.
+[34]).
+This behaviour changes, when considering the case with two outer layers that is represented by the red curve in Figure 10.
+In these models the onset for dynamo action is shifted to larger magnetic Reynolds numbers so that the minimum is now which is found for the flow field obtained at .
+Furthermore, the range of flow fields with a (relatively) small critical magnetic Reynolds number narrows and only spans the interval .
+Interestingly, although we still use the same structure for the flow fields, we get a large even when the double roll mode is present with a significant amplitude.
+This indicates that the impact of the double roll mode might be less significant than expected from the cases without outer layers but that the electromagnetic coupling of the field induced essentially by shear close to the fluid boundary layers with the side walls with finite electrical conductivity has a similarly large influence.
+IV Full set of magnetohydrodynamic equations
+In the following we present results obtained from the simulation of the induction equation coupled with the Navier-Stokes equation in a self-consistent MHD model.
+IV.1 Code scheme
+The full magnetohydrodynamic problem additionally requires a computation of the temporal evolution of the flow field.
+In this case, the Navier-Stokes equation needs to be solved as well, which for a precessing flow in the rotating frame of reference takes the form
+In addition to the terms commonly encountered in regular flow problems describing nonlinear interactions and dissipation, we have three additional contributions: the Coriolis force due to rotation, the Lorentz force due to the interaction with the magnetic field, and the Poincaré force accounting for acceleration due to perpetual change of the orientation of the cylinder’s axis of rotation.
 A detailed derivation of the Navier-Stokes equation for a precessing system is given e.g.
@@
 The drawback of this approach is the requirement of periodic boundary conditions in all three directions.
-Since this rules out an application of a cylindrical geometry, the scheme is complemented by a direct forcing immersed boundary method (IBM) which allows the consideration of nearly arbitrary shaped boundaries that prescribe the confinement for the fluid flow when the boundary conditions on these boundaries are of no-slip type.
-The code is parallelized using a slab decomposition and a standard MPI library.
-Since the most important part of the calculations results from the computation of the Fast Fourier Transformation (FFT), special efforts are undertaken to apply an extremely efficient scheme The fulfilment of the divergence-free condition for the velocity field is ensured using a standard Pressure-Corrector method where the pressure serves as a Lagrangian multiplier so that is ensured (i.e.
-first, solve Equation (2) without pressure, then project the intermediate solution on the divergence-free space by calculating the necessary pressure field).
-The same scheme is also applied to the magnetic field to ensure that the solution of the induction equation is also divergence-free as required by Maxwell’s equations.
-The code has been tested with various configurations that proved accurate solutions with excellent convergence properties and nearly perfect strong scaling properties.
-A detailed description is given in Wilbert, Giesecke, and Grauer [14] and a comparable approach was used for numerical models of the VKS dynamo[32, 33].
-III.2 Model set-up
-We consider a set-up given by a cylindrical fluid region with radius and height embedded in a Cartesian domain with a edge length so that the size of the box provides sufficient space to let the magnetic field decline outside the cylindrical vessel.
-The cylindrical vessel is implemented by imposing internal conditions on the fluid that emulate the (virtual) container walls by enforcing no-slip conditions at the surface of the container applying the direct forcing immersed boundary method and we do not model any special region representing the wall of the vessel because of such a layer would not be properly resolved.
-Instead we use a magnetic diffusivity for the fluid region, whereas the ’non-liquid’ region outside the vessel has a magnetic diffusivity , which is supposed to emulate an electrically isolating exteriour.
-The simulations are carried out with a Reynolds number of which is the maximum possible at the resolution of that still sufficiently resolves the boundary layers at the fluid-wall interface of the vessel [31].
-III.3 Results
-Similar to the procedure in the kinematic case, the Poincaré number is varied, whereby different magnetic Reynolds numbers are considered, beginning with high values.
+Since this rules out an application of a closed cylindrical geometry, the scheme is complemented by a direct forcing immersed boundary method (IBM) which allows the consideration of nearly arbitrary shaped boundaries that prescribe the confinement for the fluid flow when the boundary conditions on these boundaries are of no-slip type.
+The fulfilment of the divergence-free condition for the velocity field is ensured using a Pressure-Corrector method.
+The same scheme is also applied to the magnetic field to ensure that the solution of the induction equation is divergence-free as required by Maxwell’s equations.
+The code is parallelized using a two-dimensional processor grid.
+To apply the Fast Fourier Transform (FFT) to the distributed data along all spatial directions, the processor grid must be transposed, which is achieved by leveraging advanced MPI functions to perform an all-to-all communication of the processor data.
+Tests have been conducted with various configurations that proved accurate solutions with excellent convergence properties and nearly perfect strong scaling properties.
+A detailed description is given in Wilbert, Giesecke, and Grauer [13] and in Wilbert [44] and a comparable approach was used for numerical models of the VKS dynamo[45, 46].
+IV.2 Model set-up
+We consider a setup comparable to the kinematic models with a stagnant outer layer, i.e.
+a setup given by a cylindrical fluid region with radius and height embedded in a Cartesian domain with a edge length so that the size of the box provides sufficient space to let the magnetic field decline outside the cylindrical vessel.
+The cylindrical vessel is implemented by imposing internal conditions on the fluid that emulate the (virtual) container walls by enforcing no-slip conditions at the surface of the container.
+Unlike in the kinematic (multi-)layer models we do not model any special region representing the wall of the vessel because such a layer would not be properly resolved.
+Instead we use a magnetic diffusivity for the fluid region, whereas the ’non-liquid’ region outside the vessel has a magnetic diffusivity , which is supposed to emulate an electrically isolating exterior.
+All simulations are carried out with a Reynolds number which is the maximum possible at the resolution of while still sufficiently resolving the boundary layers at the fluid-wall interface of the vessel [44].
+IV.3 Results
+Similar to the procedure in the kinematic case, we vary the Poincaré number , whereby different magnetic Reynolds numbers are considered, beginning with high values.
 If a positive mean growth rate is found for a particular set of parameters, further runs are executed consecutively lowering the value of until no more dynamo action can be observed.
-Rather similar to the kinematic models we find a limited range for the preferred occurrence of a dynamo in the interval .
-However, the critical magnetic Reynolds number is significantly higher with a minimum value of for .
-The slightly larger value for the optimum compared to the kinematic models can be explained by the (slightly) smaller (see also [15]).
-In the following we focus on the particular case and and compare the properties of this DNS solution with the kinematic approximation obtained at and .
-This paradigmatic examples well characterize the solution and allow a faster growth with shorter time period until a statistically steady state, while showing similar qualitative behavior as the dynamos found very close to the threshold.
-Here we focus on this particular and compare the properties of this DNS solution with and in the kinematic approximation.
-III.3.1 Temporal evolution of the kinetic energy
-The temporal evolutions of kinetic and magnetic energy are displayed in Fig. 6, where the kinetic and the magnetic energy are defined as
-The simulations start with the pure hydrodynamic case and the effects by magnetic induction are switched on at .
-After a short initial phase the magnetic energy begins to grow for about cylinder rotations and finally goes into saturation.
-The induced magnetic energy is much smaller than and both energies are dominated by the contributions from the components.
+Similar to the kinematic models we find a limited range for the preferred occurrence of a dynamo in the interval .
+However, the critical magnetic Reynolds number is significantly higher with a minimum value of at .
+The slightly larger value for the optimum compared to the kinematic models can be explained by the (slightly) smaller (see also Giesecke et al.
+[14]).
+In the following we focus on the particular case and and compare the properties of this DNS solution with the results found in the kinematic case.
+The chosen paradigmatic example well characterizes the DNS solutions found very close to the threshold while allowing a faster growth with shorter time period until a statistically steady state is reached.
+IV.3.1 Temporal evolution of the kinetic energy
+The temporal evolutions of kinetic and magnetic energy and are displayed in Fig. 11, where and are defined as
+and where denotes the edge length of the cubic domain.
+The simulations start with the pure hydrodynamic case, which is processed until a quasi-stationary statistically steady state is reached.
+A detailed description of the hydrodynamic behavior is given in Wilbert, Giesecke, and Grauer [13].
+Here we focus on the properties of the magnetic field in the dynamo case and the resulting feedback on the velocity field.
+After switching on the magnetic induction at , the magnetic energy begins to grow for about cylinder rotations and finally goes into saturation.
+Both, kinetic and magnetic energy are dominated by contributions of the azimuthal components of the related field.
+However, the dominance for the magnetic field is not particularly pronounced and the induced magnetic energy remains much smaller than .
 We find that, even in the saturated state, the presence of the magnetic field does not substantially alter the structure and the amplitude of the fluid flow via the Lorentz force.
-A striking property is that unlike the kinetic energy, which remains relatively steady, the magnetic energy is characterized by quasi-periodic peaks where its value suddenly rises to times above the average.
-Although there are no substantial changes in the flow field due to the back-reaction of the magnetic field, the bursts can be clearly assigned to corresponding peaks in the interaction parameter , which compares the amplitude of the Lorentz force to the inertial term in the Navier-Stokes equation:
-which in case of a dynamo follows Miralles, Plihon, and Pinton [34].
-III.3.2 Dynamo bursts and field structure
-The peaks in the magneitc energy are always associated with a very strong concentration of magnetic energy in the form of local patches, as can be seen in the three-dimensional representation (see Fig 8).
+A striking property of the magnetic energy is that, unlike the kinetic energy, which remains relatively steady, the magnetic energy is characterized by strong occasional peaks where suddenly rises to times above its average value.
+There are no related substantial changes in the flow field due to the back-reaction of the magnetic field so that the bursts can be clearly assigned to corresponding peaks in the interaction parameter , which compares the amplitude of the Lorentz force to the inertial term in the Navier-Stokes equation:
+For a self-excited dynamo follows (see Miralles, Plihon, and Pinton [48]), which is indeed the case as shown in Figure 12, where we see similar irregular peaks due to the bursts shown by the magnetic energy in Figure 11b.
+However, the impact of the magnetic field remains small even in the case of a burst.
+In the following, we discuss three characteristic periods as labeled by capital letters A, B, (two paradigmatic peaks) and C (quiescent period) in Figure 12.
+Figure 13 shows three rows each with four snapshots of the spatial distribution of the magnetic energy taken at four arbitrary times within each related period.
+Animated sequences showing the time development within the marked periods can be found in the supplementary material of this study[47].
+We recognize a small scale field formed by flux patches distributed randomly in space and time.
 These patches are short-lived despite the strong concentration of magnetic energy, and even if they appear regularly, no corresponding period can be assigned.
-The field structure obtained from the simulations is visualized in Figure 8 which shows nested isosurfaces of the magnetic energy density for a couple of typical snapshots.
-We find a small scale field in terms of seemingly randomly distributed in space and time.
-The magnetic field patches are slightly elongated in the azimuthal direction (which goes along with the decrease of the scale in the perpendicular direction) and in general we find a spatial scale of the order of the energy carrying turbulence.
-The magnetic field patches exhibit an intermittent character and the thickness of the structures at least in one dimension is determined by the resistive scale.
-Due to the special structure of the Poincare force and the associated mean flow in the form of a standing forced inertial wave, we are not dealing with an isotropic, homogeneous turbulence, so that Kazantsev’s theory (see e.g.
-[brandenburgXY]) is not directly applicable.
-Let us label the second (major) peak as case A, the third peak as case B, and the non-peak solution as case C (see Fig.
-6).
-The magnetic field is small-scale in all cases, whereas the velocity field is large-scale.
-Although our self-exciting dynamo is small-scale, ongoing hydromagnetic processes tend to excite larger scales (i.e., a large-scale magnetic field), as seen in Figure 8a.
-and 8b.
-However, these large-scale magnetic field structures are quite isolated; they do not extend throughout the entire volume and quickly dissipate due to strong magnetic diffusion, typically within 1-2 rotation periods.
-In contrast, the velocity field is large-scale, characterized by two vortex structures near the cylinder wall, and is concentrated mainly near the cylinder boundary, while it is almost zero in the central volume due to the dominance of the component.
-The vorticity, , naturally follows the velocity characteristics, being concentrated at the cylinder’s border and negligible toward the central volume, as shown in Fig.
-LABEL:B-U-O.
-The velocity field remains largely unchanged in cases A and B, retaining characteristics similar to those in case C.
-The vorticity, however, penetrates slightly deeper into the central volume compared to case C (see Fig.
-LABEL:B-U-O), which represents the standard off-peak solution, where the magnetic field is weak and small-scale without any significant large-scale structures.
+The magnetic field patches are slightly elongated in the azimuthal direction and during a typical peak period, we see an increased excitation of medium-scale concentrations of magnetic energy in terms of elongated arc-like magnetic structures near the end caps (Figure 13a–13h).
+However, these mid-scale magnetic field structures remain isolated and decay within 1-2 rotation periods, whereby the resulting small scale patches have a rather isotropic character and are distributed within the entire volume of the cylinder.
 While the distribution of the magnetic energy in individual snapshots appears to be largely chaotic (with the exception of the bursts), this is no longer the case when we look at the time average in the turntable system.
-With sufficiently long averaging of more than 200 rotation periods, we get the distribution shown in Figure 9.
-III.3.3 Backreaction and kinetic power spectrum
-A reduction in the speed amplitude is also not absolutely necessary, because the dynamo that sets in is a small-scale dynamo.
-For this case, theoretical considerations have shown that (in a mean-field model) the feedback effect via the Lorentz force corresponds to an increased diffusivity or hyperdiffusivity [35].
-However, an reaction of the fluid flow can be seen in a subtle way when looking at the temporal power-spectra of kinetic energy.
-Figure 10 compares the temporal Power-spectrum of the kinetic energy for the hydrodynamic case (left panel) with the spectrum for the dynamo case (right panel).
-We see that, in the hydrodynamic case, the contributions to high frequencies by the axial component (spectral index ) drop significantly faster than the contributions of the horizontal components and (spectral index ).
+Applying a long averaging time of more than 200 rotation periods, we obtain a rather regular distribution of magnetic energy as shown in Figure 14.
+Similar to the low-efficient branch of the kinematic models we find elongated structures close to the end caps, which are a result of the statistical distribution of the magnetic field energy.
+Furthermore, and in contrast to the kinematic models, we recognize different symmetries with respect to the mid-plane of the cylinder, which are expressed in the different signs for the time-averaged radial field.
+We therefore see both, a dipole-like state and a quadrupole-like state, which are alternately dominant at irregular intervals.
+Due to the small-scale structure, however, the characterization as a dipole or quadrupole field is not really justified.
+In this sense, the results are consistent with other dynamo simulations in which a mechanical forcing was utilized as driving mechanism (see e.g.
+Lin et al.
+[49], Cébron et al.
+[5], Reddy, Favier, and Le Bars [50]).
+In agreement with Landeau’s conclusions[9], we therefore confirm that it remains unclear whether mechanical forcing can generate a large-scale magnetic field, and apart from experimental investigation, only simulations with massively reduced viscosity (i.e.
+massively larger Reynolds number) can help to clarify this question, which is currently prevented by insufficient performance of numerical algorithms and HPC systems.
+IV.3.2 Back-reaction and kinetic power spectrum
+In contrast to the magnetic field, the velocity field is always dominated by large scales as shown in Figure 15 by means of a paradigmatic time-snapshot of the three-dimensional distribution of the axial velocity component .
+This figure shows the typical structure of the velocity, which looks very similar at any time, regardless of whether with or without a magnetic field, or whether within a peak period or within a magnetically quiet section.
+Indeed, a reduction of the speed amplitude and/or of the field structure is not absolutely necessary, because here we deal with a small-scale dynamo for which theoretical considerations (in a mean-field model) have shown that the feedback via the Lorentz force can be described by an increased diffusivity or hyperdiffusivity (Subramanian [51]).
+A response of the fluid flow can be seen in a subtle way when looking at the temporal power-spectra of kinetic energy.
+Figure 16 compares the temporal power spectrum of the kinetic energy for the hydrodynamic case (left panel) with the spectrum for the dynamo case (right panel).
+Here we assume that in the high frequency range the kinetic energy of the individual components and scales according where is the energy from the Fourier component with the frequency and is a spectral index that turns out to be different for the horizontal coordinates () and the vertical coordinate ().
+In the hydrodynamic case, the contributions to high frequencies by the axial component (spectral index ) drop significantly faster than the contributions of the horizontal components and (spectral index ).
 In the dynamo case, the spectral index of the horizontal components approaches which is close to the axial component, which practically does not change compared to the hydrodynamic case.
-This change is not particularly strong, but it is still clearly noticeable.
-IV Conclusions
-The kinematic models reveal the existence of two different branches for dynamo action.
-One branch is characterized by a highly localized and stationary dynamo that only occurs at rather large magnetic Reynolds number.
+In summary, in the magnetic case we find an isotropization of the spectrum and a damping of the temporal fluctuations, which is expressed in the steeper spectrum.
+V Summary and Conclusions
+We have run numerical simulations of precession driven dynamo action in a cylindrical cavity.
+We examined simplified kinematic models as well as a particular setup of a similar system solved by a self-consistent three-dimensional magnetohydrodynamic approach.
+The kinematic models reveal the existence of two different branches for dynamo action, the realization of which depends on the configuration of the outer boundary regions.
+The first branch is characterized by a highly localized and stationary dynamo that only occurs at rather large magnetic Reynolds number.
 The second type of dynamo is of oscillatory type with the magnetic field filling a larger proportion of the fluid domain.
 This type can occur at much lower magnetic Reynolds number, which, however requires suitable boundary conditions in terms of either a thin outer layer with large magnetic diffusivity or a comparable emulation with pseudo-vacuum BC.
-However, the kinematic model, which we assume best reflects the experimental setup, has a critical magnetic Reynolds number of , which is about twice as high as achievable in the planned precession dynamo experiment.
-Interestingly, this agrees quite well with the value derived by Goepfert and Tilgner [20] by means of simulations in a precessing cube.
-We then applied direct numerical simulations and have calculated a comparative solution in order to investigate the time evolution of the velocity field and how it is influenced by the magnetic field through the Lorentz force and to check for consistentency with the kinematic models.
-The configuration of this DNS essentially corresponds to that of the kinematic model, in which a stationary outer layer with large magnetic diffusivity emulates the outer space in the laboratory.
-Unfortunately, it is not possible to achieve the large Reynolds numbers, as they are relevant in the experiment, due to the limitations of the numerics and the hardware, respectively.
-This is mainly due to the extremely thin boundary layers, which require a sufficiently high resolution.
-As a consequence, we find dynamo solutions in the DNS only for relatively large magnetic Reynolds numbers so that we effectively deal with a magnetic Prandtl number of the order of one with the magnetic field being small-scale and fluctuating on short time scales.
-At this stage it cannot be concluded with certainty that this behavior will change if we move to larger Reynolds numbers, as there are currently insufficient parameter studies available for such scaling.
-For the parameters considered, a small-scale burst dynamo is the typical solution.
-These bursts are characterized by peaks in magnetic energy, which, as previously mentioned, represent the sudden excitation of large-scale magnetic field structures.
-Since these structures are isolated and not sustained throughout the entire volume, they are rapidly destroyed by strong magnetic diffusion.
-This is followed by a period without the excitation of large-scale structures (no peaks in magnetic energy) until the hydromagnetic dynamo once again excites some large-scale magnetic field structures.
-Our results align with findings [36, 37, 38], which conclude that increasing the magnetic Prandtl number leads to a transition from large-scale dynamos to small-scale dynamos.
-Thus, at high values of , small-scale dynamos predominate, which is consistent with our case.
-To summarize, the characteristics of the magnetic fields from the kinematic simulations are clearly different from those of the DNS.
+We have not investigated here at which exact parameters the transition from one branch to the other takes place, since the experimental realization must in any case start from a basically infinitely large outer volume of an insulator, namely the air in the laboratory.
+The kinematic model with two outer layers (for the container walls and for the lab environment), which we assume best reflects the experimental setup, has a critical magnetic Reynolds number of , which agrees surprisingly well with the value derived by Goepfert and Tilgner [30] obtained from simulations in a precessing cube.
+Our solutions with external outer walls and an increased diffusivity (i.e.
+a decreased electrical conductivity) are consistent with previous studies[52, 53], who found a monotonic reduction of the dynamo threshold for increasing conductivity of an outer wall.
+Furthermore, these studies also reported a reduction for oscillatory solutions with an optimum at a particular thickness of the outer wall, which was explained by the increased dissipation resulting from eddy currents induced in the outer wall.
+We also used direct numerical simulations in order to investigate the time evolution of the velocity field, the impact of the magnetic field through the Lorentz force and the consistency with the kinematic models.
+Unfortunately, it is not possible to achieve the large hydrodynamic Reynolds number, as it is relevant in the experiment, due to the limitations of the numerics and the hardware, so that we find dynamo solutions in the DNS only for relatively large magnetic Reynolds numbers.
+This means that we effectively deal with a magnetic Prandtl number of the order of one and that the magnetic field is of small-scale type.
+At this stage it cannot be concluded with certainty that this behavior will change if we move to larger (hydrodynamic) Reynolds numbers, as there are currently insufficient parameter studies available for such scaling.
+Nevertheless, for the parameters considered, a small-scale burst dynamo is the typical solution.
+These bursts are characterized by peaks in magnetic energy, which represent the sudden excitation of large-scale magnetic field structures.
+Since these structures are isolated and not sustained throughout the entire volume, they are rapidly destroyed by magnetic diffusion.
+Our results align with findings obtained for example by Etchevest, Fontana, and Dmitruk [54], Richardson, Hollerbach, and Proctor [55], Hughes, Cattaneo, and Thelen [56], which conclude that increasing the magnetic Prandtl number leads to a transition from large-scale dynamos to small-scale dynamos.
+The small scale characteristics of the magnetic fields from the DNS are clearly different from the properties of the field obtained in the kinematic simulations.
 However, if the magnetic field from the DNS is averaged over a sufficiently long period of time so that the occasional occurrence of the strongly localized patches is also averaged out, the structure is comparable to the structure of the weakly efficient branch in the kinematic models, which only appears at large magnetic Reynolds numbers.
 This explains, at least in part, the high magnetic Reynolds number required to obtain a dynamo in the DNS.
-Averaging the magnetic fields over a sufficient period of time, the small scale magnetic field structures give a similar pattern as for the weak branch of the kinematic solutions, while the occasional bursts appear to have a similar character to the strong branch, but fail to establish a corresponding dynamo solution.
+In order to obtain a dynamo solution that can also be realized in the planned experiment, it would be necessary to find a path that leads to the efficient dynamo branch with low .
+Whether this exists at all in the full MHD case, and how this can be achieved, remains unknown at present.
 Acknowledgements.
@@
 (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC).
-Ján Šimkanin is grateful for support from the Alexander-von-Humboldt Stiftung (CZE 1079936).References
+Ján Šimkanin is grateful for support from the Alexander-von-Humboldt Stiftung (CZE 1079936) and for funding from the Helmholtz Association in frame of the AI project GEOMAGFOR (ZT-I-PF-5-200).References
 - Stieglitz and Müller [2001] R. Stieglitz and U. Müller, “Experimental demonstration of a homogeneous two-scale dynamo,” Phys. Fluids 13, 561–564 (2001).
@@
 98, 044502 (2007), 10.1103/PhysRevLett.98.044502.
+- Stefani et al.
+[2015] F.
+Stefani, T.
+Albrecht, G.
+Gerbeth, A.
+Giesecke, T.
+Gundrum, J.
+Herault, C.
+Nore, and C.
+Steglich, “Towards a precession driven dynamo experiment,” Magnetohydrodynamics 51, 275–284 (2015).
+- Cébron et al. [2019] D. Cébron, R. Laguerre, J. Noir, and N. Schaeffer, “Precessing spherical shells: flows, dissipation, dynamo and the lunar core,” Geophys. J. Int. 219, S34–S57 (2019).
+- Malkus [1968] W. V. R. Malkus, “Precession of the Earth as the Cause of Geomagnetism,” Science 160, 259–264 (1968).
+- Stys and Dumberry [2020] C. Stys and M. Dumberry, “A past Lunar dynamo thermally driven by the precession of its inner core,” Journal of Geophysical Research: Planets 125, e2020JE006396 (2020).
+- Olson [2013] P. Olson, “The new core paradox,” Science 431–432, 013008 (2013).
+- Landeau et al. [2022] M. Landeau, A. Fournier, H.-C. Nataf, D. Cébron, and N. Schaeffer, “Sustaining earth’s magnetic dynamo,” Nature Reviews Earth and Environment 3, 255 – 269 (2022).
+- Rochester et al. [1975] M. G. Rochester, J. A. Jacobs, D. E. Smylie, and K. F. Chong, “Can precession power the geomagnetic dynamo?” Geophys. Journal 43, 661–678 (1975).
+- Kerswell [1996] R. R. Kerswell, “Upper bounds on the energy dissipation in turbulent precession,” J. Fluids Mech. 321, 335–370 (1996).
+- Pizzi et al. [2021] F. Pizzi, A. Giesecke, J. Šimkanin, and F. Stefani, “Prograde and retrograde precession of a fluid -filled cylinder,” New J. Phys. 23, 123016 (2021).
+[…]
+- Giesecke et al. [2024] A. Giesecke, T. Vogt, F. Pizzi, V. Kumar, F. Garcia-Gonzalez, T. Gundrum, and F. Stefani, “The global flow state in a precessing cylinder.” J. Fluid Mech. 998, A30 (2024).
+- Kobine [1996] J. J. Kobine, “Azimuthal flow associated with inertial wave resonance in a precessing cylinder,” J. Fluid Mech. 319, 387 – 406 (1996).
+- Manasseh [1992] R. Manasseh, “Breakdown regimes of inertia waves in a precessing cylinder,” J. Fluids Mech. 243, 261–296 (1992).
+- Manasseh [1994] R. Manasseh, “Distortions of inertia waves in a rotating fluid cylinder forced near its fundamental mode resonance,” J. Fluids Mech. 265, 345–370 (1994).
+- Manasseh [1996] R. Manasseh, “Nonlinear behaviour of contained inertia waves,” J. Fluids Mech. 315, 151–173 (1996).
+- Noir et al. [2001] J. Noir, D. Brito, K. Aldridge, and P. Cardin, “Experimental evidence of inertial waves in a precessing spheroidal cavity,” Geophys. Res. Lett. 28, 3785–3788 (2001).
+- Noir et al.
+[2003] J.
+Noir, P.
+Cardin, D.
+Jault, and J.-P.
+Masson, “Experimental evidence of non-linear resonance effects between retrograde precession and the tilt-over mode within a spheroid,” Geophys.
+J.
+Int.
+154, 407–416 (2003).
+- Mouhali et al. [2012] W. Mouhali, T. Lehner, J. Léorat, and R. Vitry, “Evidence for a cyclonic regime in a precessing cylindrical container,” Exp. Fluids 53, 1693–1700 (2012).
+[…]
+- Le Bars et al.
+[2022] M.
+Le Bars, A.
+Barik, F.
+Burmann, D.
+P.
+Lathrop, J.
+Noir, N.
+Schaeffer, and S.
+A.
+Triana, “Fluid dynamics experiments for planetary interiors,” Surveys in Geophysics 43, 229 – 261 (2022).
+- Burmann, Kira, and Noir [2024] F.
+Burmann, L.
+Kira, and J.
+Noir, “Precessing non-axisymmetric ellipsoids: bi-stability and fluid instabilities,” Journal of Fluid Mechanics 997 (2024), 10.1017/jfm.2024.774.
+- Léorat [2006] J. Léorat, “Large scales features of a flow driven by precession,” Magnetohydrodynamics 42, 143–151 (2006).
+- Christensen and Aubert [2006] U.
+R.
+Christensen and J.
+Aubert, “Scaling properties of convection-driven dynamos in rotating spherical shells and application to planetary magnetic fields,” Geophysical Journal International 166, 97–114 (2006).
+- King et al.
+[2010] E.
+M.
+King, K.
+M.
+Soderlund, U.
+R.
+Christensen, J.
+Wicht, and J.
+M.
+Aurnou, “Convective heat transfer in planetary dynamo models,” Geochemistry, Geophysics, Geosystems 11 (2010), https://doi.org/10.1029/2010GC003053.
+[…]
+- Wu and Roberts [2009] C.-C. Wu and P. Roberts, “On a dynamo driven by topographic precession,” Geophys. Astrophys. Fluid Dyn. 103, 467–501 (2009).
+- Goepfert and Tilgner [2016] O. Goepfert and A. Tilgner, “Dynamos in precessing cubes,” New J. Phys. 18, 103019 (2016), 10.1088/1367-2630/18/10/103019.
+- Nore et al. [2011] C. Nore, J. Léorat, J.-L. Guermond, and F. Luddens, “Nonlinear dynamo action in a precessing cylindrical container,” Phys. Rev. E 84, 016317 (2011).
+- Cappanera et al. [2016] L. Cappanera, J.-L. Guermond, J. Léorat, and C. Nore, “Two spinning ways for precession dynamo,” Phys. Rev. E 93, 043113 (2016).
 - Giesecke et al.
@@
 120, 024502 (2018).
-- Stefani et al.
-[2019] F.
-Stefani, A.
-Gailitis, G.
-Gerbeth, A.
-Giesecke, T.
-Gundrum, G.
-Rüdiger, M.
-Seilmayer, and T.
-Vogt, “The dresdyn project: liquid metal experiments on dynamo action and magnetorotational instability,” Geophys.
-Astrophys.
-Fluid Dyn.
-113, 51 – 70 (2019).
-- Léorat [2006] J. Léorat, “Large scales features of a flow driven by precession,” Magnetohydrodynamics 42, 143–151 (2006).
-- Malkus [1968] W. V. R. Malkus, “Precession of the Earth as the Cause of Geomagnetism,” Science 160, 259–264 (1968).
-- Cébron et al. [2019] D. Cébron, R. Laguerre, J. Noir, and N. Schaeffer, “Precessing spherical shells: flows, dissipation, dynamo and the lunar core,” Geophys. J. Int. 219, S34–S57 (2019).
-- Olson [2013] P. Olson, “The new core paradox,” Science 431–432, 013008 (2013).
-- Landeau et al. [2022] M. Landeau, A. Fournier, H.-C. Nataf, D. Cébron, and N. Schaeffer, “Sustaining earth’s magnetic dynamo,” Nature Reviews Earth and Environment 3, 255 – 269 (2022).
-- Rochester et al. [1975] M. G. Rochester, J. A. Jacobs, D. E. Smylie, and K. F. Chong, “Can precession power the geomagnetic dynamo?” Geophys. Journal 43, 661–678 (1975).
-- Kerswell [1996] R. R. Kerswell, “Upper bounds on the energy dissipation in turbulent precession,” J. Fluids Mech. 321, 335–370 (1996).
-- Pizzi et al. [2021] F. Pizzi, A. Giesecke, J. Šimkanin, and F. Stefani, “Prograde and retrograde precession of a fluid -filled cylinder,” New J. Phys. 23, 123016 (2021).
-[…]
-- Giesecke et al.
-[2024] A.
-Giesecke, T.
-Vogt, F.
-Pizzi, V.
-Kumar, F.
-Garcia-Gonzalez, T.
-Gundrum, and F.
-Stefani, “The global flow state in a precessing cylinder.” J.
-Fluid Mech.
-XXX, in press (2024).
-- Stys and Dumberry [2020] C. Stys and M. Dumberry, “A past Lunar dynamo thermally driven by the precession of its inner core,” Journal of Geophysical Research: Planets 125, e2020JE006396 (2020).
-[…]
-- Wu and Roberts [2009] C.-C. Wu and P. Roberts, “On a dynamo driven by topographic precession,” Geophys. Astrophys. Fluid Dyn. 103, 467–501 (2009).
-- Nore et al. [2011] C. Nore, J. Léorat, J.-L. Guermond, and F. Luddens, “Nonlinear dynamo action in a precessing cylindrical container,” Phys. Rev. E 84, 016317 (2011).
-- Goepfert and Tilgner [2016] O. Goepfert and A. Tilgner, “Dynamos in precessing cubes,” New J. Phys. 18, 103019 (2016), 10.1088/1367-2630/18/10/103019.
-- Gans [1971] R. F. Gans, “On hydromagnetic precession in a cylinder,” J. Fluids Mech. 45, 111–130 (1971).
 - Giesecke et al.
@@
 113, 235–255 (2019).
-- Pizzi, Giesecke, and Stefani [2021] F. Pizzi, A. Giesecke, and F. Stefani, “Ekman boundary layers in a fluid filled precessing cylinder,” AIP Advances 11, 035023 (2021).
-- Blackburn and Sherwin [2004] H.
-M.
-Blackburn and S.
-J.
-Sherwin, “Formulation of a Galerkin spectral element-Fourier method for three-dimensional incompressible flows in cylindrical geometries,” J.
-Comp.
-Phys.
-197, 759–778 (2004).
-- Blackburn et al.
-[2019] H.
-M.
-Blackburn, D.
-Lee, T.
-Albrecht, and J.
-Singh, “Semtex: a spectral element–Fourier solver for the incompressible Navier–Stokes equations in cylindrical or Cartesian coordinates,” Comput.
-Phys.
-Cmm.
-245, 106804–1–13 (2019).
+- Pizzi et al.
+[2022] F.
+Pizzi, A.
+Giesecke, J.
+Šimkanin, V.
+Kumar, T.
+Gundrum, and F.
+Stefani, “Numerical and theoretical framework for the dresdyn precession dynamo experiment,” Magnetohydrodynamics 58, 445 – 453 (2022).
+- Kumar et al.
+[2023] V.
+Kumar, F.
+Pizzi, A.
+Giesecke, J.
+Šimkanin, T.
+Gundrum, M.
+Ratajczak, and F.
+Stefani, “The effect of nutation angle on the flow inside a precessing cylinder and its dynamo action,” Phys.
+Fluids 35 (2023), 10.1063/5.0134562.
 - Giesecke, Stefani, and Gerbeth [2008] A.
@@
 Gerbeth, “Kinematic simulation of dynamo action by a hybrid boundary-element/finite-volume method,” Magnetohydrodynamics , 237 – 252 (2008).
+- Guermond et al.
+[2009] J.-L.
+Guermond, R.
+Laguerre, J.
+Léorat, and C.
+Nore, “Nonlinear magnetohydrodynamics in axisymmetric heterogeneous domains using a fourier/finite element technique and an interior penalty method,” Journal of Computational Physics 228, 2739 – 2757 (2009).
 - Giesecke et al.
@@
 Luddens, and J.-L.
-Guermondz, “Electromagnetic induction in non-uniform domains,” Geophys.
+Guermond, “Electromagnetic induction in non-uniform domains,” Geophys.
 Astrophys.
@@
 104, 505 – 529 (2010).
-- Giesecke [2024a] A. Giesecke, “supplemental material,” www.giesecke.net (2024a).
-- Giesecke [2024b] A. Giesecke, “supplemental material,” www.giesecke.net (2024b).
-- Rüdiger, Elstner, and Ossendrijver [2003] G. Rüdiger, D. Elstner, and M. Ossendrijver, “Do spherical -dynamos oscillate?” Astron. Astrophys. 406, 15–21 (2003).
+- Rädler et al.
+[2002] K.-H.
+Rädler, M.
+Rheinhardt, E.
+Apstein, and H.
+Fuchs, “On the mean-field theory of the karlsruhe dynamo experiment i.
+kinematic theory,” Magnetohydrodynamics 38, 41 – 71 (2002).
+- Pizzi, Giesecke, and Stefani [2021] F. Pizzi, A. Giesecke, and F. Stefani, “Ekman boundary layers in a fluid filled precessing cylinder,” AIP Advances 11, 035023 (2021).
+- Dudley and James [1989] M. L. Dudley and R. W. James, “Time-dependent kinematic dynamos with stationary flows,” Proceedings of the Royal Society of London Series A 425, 407–429 (1989).
+- Giesecke [2024a] A. Giesecke, “supplemental material,” will be inserted by publisher (2024a).
 - Wilbert [2024] M.
@@
 119, 234501 (2017).
+- Giesecke [2024b] A. Giesecke, “supplemental material,” will be inserted by publisher (2024b).
 - Miralles, Plihon, and Pinton [2015] S.
@@
 775, 501–523 (2015).
+[…]
+- Reddy, Favier, and Le Bars [2018] K.
+S.
+Reddy, B.
+Favier, and M.
+Le Bars, “Turbulent kinematic dynamos in ellipsoids driven by mechanical forcing,” Geophysical Research Letters 45, 1741 – 1750 (2018).
 - Subramanian [2003] K. Subramanian, “Hyperdiffusion in nonlinear large- and small-scale turbulent dynamos,” Phys. Rev. Lett. 90, 245003 (2003).
+- Avalos-Zuniga, Plunian, and Gailitis [2003] R.
+Avalos-Zuniga, F.
+Plunian, and A.
+Gailitis, “Influence of electromagnetic boundary conditions onto the onset of dynamo action in laboratory experiments,” Phys.
+Rev.
+E 68, 0663071 (2003).
+- Avalos-Zuñiga and Plunian [2005] R.
+Avalos-Zuñiga and F.
+Plunian, “Influence of inner and outer walls electromagnetic properties on the onset of a stationary dynamo,” European Physical Journal B 47, 127–135 (2005).
 - Etchevest, Fontana, and Dmitruk [2022] M.